### v4

# A variational formulation of the free energy of mixed quantum-classical systems: coupling classical and electronic density functional theories

Abstract
Combining classical density functional theory (cDFT) with quantum mechanics (QM) methods offers a computationally efficient alternative to traditional QM/molecular mechanics (MM) approaches for modeling mixed quantum-classical systems at finite temperatures. However, both QM/MM and QM/cDFT rely on somewhat ambiguous approximations.
This paper aims to develop an exact theoretical framework that allows us to clarify the approximations involved in the QM/cDFT formulation and serves as a starting point for further development. We therefore establish a comprehensive density functional theory (DFT) framework for mixed quantum-classical systems within the canonical ensemble. We start with the adiabatic equilibrium density matrix for a system comprising $N_{{\text{qm}}}$ quantum and $N_{{\text{mm}}}$ classical particles and, based on this, we propose a variational formulation of the Helmholtz free energy in terms of the full QM/MM density matrix. Taking advantage of permutational symmetry and thanks to constrained-search methods, we reformulate the computation of the Helmholtz free energy using only the quantum and classical one-body densities, thus generalizing both cDFT and electronic DFT (eDFT) for QM/MM systems.
We then reformulate the functional in order to explicitly make appear the eDFT and cDFT Levy-Lieb functionals, together with a new universal correlation functional for QM/MM systems. A mean-field approximation is finally introduced in the context of solvation problems and we discuss its connection with several existing mixed cDFT-eDFT schemes.
I Introduction
The Hohenberg-Kohn theoremsHohenberg and Kohn (1964) dramatically reduce the complexity of simulating a quantum (QM) molecular system by enabling a focus on the one-electron density instead of the full $N$-electron wave function. Central to this simplification is the unknown universal density functional, which was later given a more rigorous mathematical foundation by Levy and LiebLevy (1979); Lieb (1983). Through the self-consistent framework introduced by Kohn and ShamKohn and Sham (1965), along with extensive efforts to develop approximations to the exact exchange-correlation functional (see, for example, a recent review in Ref. Toulouse, 2022), electronic density functional theory (eDFT) has become the workhorse of quantum chemistry.
Although electronic structure calculations at zero temperature can routinely handle systems with hundreds to even thousands of atoms, accounting for finite-temperature effects becomes crucial in many applications. Such calculations must therefore be set in a statistical ensemble, where the leading-order entropic contribution arises from nuclear motion. In this context, it is natural to compute the forces acting on the nuclei at the quantum level and then to solve the classical equations of motion to propagate the trajectory. This forms the basis of ab-initio molecular dynamics Iftimie, Minary, and Tuckerman (2005); Silvestrelli and Parrinello (1999). Despite its considerable achievements, it is computationally demanding due to the extensive number of QM calculations required to sample the exponentially large phase space. The convergence rate of the statistics can also vary significantly depending on the targeted properties and conditionsPan et al. (2013); Rozsa et al. (2018). To alleviate ab-initio molecular dynamics computational burden, QM/MM methods were developed, where only a small, critical portion of the system is treated at the QM level, while the remainder is handled at the classical level (MM). The interaction between the classical particles are modeled using parametrized force fields and their dynamics obey Newton’s law of motion. However, even in QM/MM approaches, sampling the phase space of large systems remains computationally intensive.
Inspired by the extension of eDFT to QM systems at finite temperaturesMermin (1965), density-based formalisms have emerged for describing purely classical objects in statistical mechanics, particularly in the form of classical density functional theory (cDFT)Evans (1979). Originally rooted in liquid state theory, cDFT is essentially a classical counterpart to eDFT, reformulating the classical statistical mechanics problem of large ensembles of identical particles (atoms or molecules) in terms of one-body density functionals. cDFT, which is naturally formulated within the grand canonical ensemble, establishes the existence of a unique functional of the classical particle density that reaches its minimum at the equilibrium density. At this minimum, the functional is equal to the grand potential, which serves as the thermodynamic potential of the grand canonical ensemble. Therefore, most equilibrium properties can be obtained through functional minimization over the classical one-body density rather than through tedious sampling of configurations in an exponentially large phase space.
The efficiency of both eDFT and cDFT stems from a common principle: each method replaces the challenge of computing a high-dimensional object—such as the wavefunction in QM or the probability distribution in MM—with the optimization of a functional that depends only on a low-dimensional quantity, i.e. the one-body density of either quantum electrons or classical molecules. The mathematical framework allowing for such a simplification is the so-called constrained-search principle introduced by Levy and LiebLevy (1979); Lieb (1983) which enables, under certain constraints of representativity, to give an explicit expression of the universal functionals of eDFT and cDFT, although not usable in practical calculations.
Its computational efficiency makes cDFT appealing to model solvent effect at the classical level while retaining a quantum description of the soluteTang et al. (2020); Petrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024). In this approach, the total Hamiltonian is partitioned into well defined QM and MM Hamiltonians.However, the QM/MM interaction lacks of a unique practical definition. The coupling Hamiltonian is usually described as the sum of an electrostatic interaction combined with a non-bonded term which is often modeled using Lennard-Jones force field to account for repulsion and dispersion effectsSenn and Thiel (2009); Giovannini, Lafiosca, and Cappelli (2017); Clemente, Capece, and Martí (2023). The Mixed cDFT-eDFT approach advantageously replaces the exhaustive phase-space sampling required in QM/MM molecular dynamics with a self-consistent optimization of both classical and quantum functionals. While the idea of merging cDFT and quantum methods is completely natural its theoretical justification is not immediately straightforward.
Petrosyan and co-workersPetrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007) justified their so-called joint density-functional formulation by the use of a Levy-Lieb constrained search procedure allowing to express the universal functional as depending on the electron density and the nuclear density rather than the full density matrix describing the electrons and nuclei. To derive a QM/MM formulation, they treat the environment particles as classical by ’integrating out’ the electronic density associated with these nuclei. Howvever, this step makes it difficult to pursue a rigorous derivation and their working functional is rather defined as the sum of the Quantum Kohn-Sham functional describing the electrons of the solute, the classical functional describing the solvent and a coupling functional defined as the difference between the exact functional and the sum of the two former functionals. In other workJeanmairet, Levesque, and Borgis (2020); Tang et al. (2020), a similar form of the functional was invoked without further theoretical justification.
Thus, while QM/cDFT approaches offer potentially attractive computational scaling, their derivation from an exact theory remains somewhat ambiguous. This work addresses this gap by establishing a rigorous mathematical framework that provides a DFT-equivalent description of mixed QM/MM systems. Here, we focus on the canonical ensemble, deriving equations to compute the Helmholtz free energy using only the QM and MM one-body densities. Extending this formulation to the grand canonical ensemble for classical particles is conceptually straightforward but introduces additional complexity in the notations due to the varying number of particles. The main difference between the two ensembles is that, in the grand canonical ensemble, there exists a unique correspondence between the equilibrium density and the external potential. In contrast, in the canonical ensemble, the equilibrium density corresponds to a family of external potentials, defined up to an additive constantLutsko (2022).
Within the present paper, the foundational object from which the QM/MM DFT formulation is derived is the explicit expression of the adiabatic equilibrium density of a mixed QM/MM system in the canonical ensemble. This expression, which was previously established in the adiabatic basis by Nielsen et. al. in the context of QM/MM dynamicsNielsen, Kapral, and Ciccotti (2001), is recovered here following an alternative route (see Sec. II of the supplementary materials). Building on this, we propose a functional formulation of the free energy in terms of general QM and MM high-dimensional densities (see Sec. Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ for a proof of the variational properties). By applying the Levy-Lieb constrained search and assuming additive two-body QM/MM interactions, we then naturally express the free energy in terms of the one-electron density and the classical one-body density. We explicitly establish the connection between eDFT and cDFT through the introduction of a new universal QM/MM correlation one-body density functional. Additionally, we discuss a mean-field approximation and its relationship with previously proposed cDFT/eDFT couplings in the context of QM/MM solvation problems. We also note that this work can be seen as the QM/MM and finite-temperature counterpart of the recently published study by Fromager and LasorneFromager and Lasorne (2024), which formulated an exact DFT framework incorporating both electrons and nuclei at the quantum level—going beyond the Born-Oppenheimer approximation.
The paper is organized as follows. In Sec. II, we begin by discussing the adiabatic equilibrium density for a mixed QM/MM system in the canonical ensemble (Sec. II.1 and II.2). Then, in Sec. II.3, using the explicit form of the equilibrium density, we reformulate the computation of the free energy as a functional minimization over general high-dimensional mixed QM/MM densities. Sec. III focuses on expressing this problem explicitly in terms of the one-body QM and MM densities. In Sec. III.1, we employ the Levy-Lieb constrained search to reformulate the high-dimensional QM/MM functional problem by recognizing a many-to-one correspondence between the high-dimensional QM/MM quantities and the one-body QM and MM densities. In Sec. III.2, we establish a connection between eDFT and cDFT by introducing a universal QM/MM correlation functional of the one-body QM and MM densities. Sec. III.3 then discusses this formulation in the context of solvation problems and recent QM/cDFT coupling approaches. Finally, Sec. IV summarizes the main results and conclusions. The supplementary materials provide an alternative derivation of the equilibrium density for a QM/MM system that does not rely on Liouvillian operators, and the appendix Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ provides a proof of the variational properties of the QM/MM functional.
II Quantum-classical system in the Canonical ensemble: equilibrium density and variational formulation
The main goal of this section is to give a variational formulation for the Helmholtz free energy of a mixed QM/MM system based on a functional minimization over general densities involving $N_{{\text{qm}}}$ and $N_{{\text{mm}}}$ variables. We define in Sec. II.1 our working Hamiltonian describing the QM/MM system. The equilibrium density of a mixed QM/MM system is then given in Sec. II.2, and we briefly discuss the physical meaning of that result. We eventually show in Sec. II.3 that the Helmholtz free energy can be obtained through a variational principle over mixed QM/MM density matrices.
II.1 The Hamiltonian of the quantum-classical system
We adopt the following notation conventions : i) lowercase letters will refer to purely quantum objects (both variables and operators) while capital letters will refer to either purely classical or mixed quantum-classical objects and ii) all bold letters refer to collective variables or operators (i.e. relative to all quantum or classical particles) and non bold letters will refer to individual or pairwise variables or operators.
The total system is divided into a set of $N_{{\text{mm}}}$ identical classical particles, described by their classical positions ${\bf{Q}}=\{{{Q}}_{1},\ldots,{{Q}}_{N_{{\text{mm}}}}\}$ and momentum ${\bf P}=\{{{P}}_{1},\ldots,{{P}}_{N_{{\text{mm}}}}\}$, and a set of $N_{{\text{qm}}}$ identical quantum particles described by their position variables ${\bf{q}}=\{{{q}}_{1},\ldots,{{q}}_{N_{{\text{qm}}}}\}$.
The Hamiltonian of the total system is written as
$\hat{\bf H}_{\text{tot}}({\bf{Q}},{\bf P},{\bf{q}})=\hat{\bf h}_{{\text{qm}}}({\bf{q}})+{\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}},{\bf{q}}).$ (1)
In the rest of the paper, explicit dependency on the QM and MM variables will be dropped for concision when unambiguous. In Eq.(1), $\hat{\bf h}_{{\text{qm}}}$ is the quantum Hamiltonian operator describing $N_{{\text{qm}}}$ quantum identical particles of mass $m$ (typically electrons) in interaction with an external potential $v^{\text{ext}}_{{\text{qm}}}({{q}})$ local in position representation, and interacting among themselves with a pairwise potential ${w}_{{\text{qm}}}({{q}}_{1},{{q}}_{2})$. The explicit form of $\hat{\bf h}_{{\text{qm}}}$ reads then
$\hat{\bf h}_{{\text{qm}}}=\hat{\bf t}+\hat{{\bf v}}^{\text{ext}}_{{\text{qm}}}+\hat{\bf w}_{{\text{qm}}},$ (2)
with
$\hat{\bf t}=\sum_{i=1}^{N_{{\text{qm}}}}\frac{\hat{p}_{i}^{2}}{2m},\qquad\hat{p}_{i}=-i\hbar\nabla_{{{q}}_{i}},$ (3)
$\hat{{\bf v}}^{\text{ext}}_{{\text{qm}}}=\sum_{j=1}^{N_{{\text{qm}}}}v^{\text{ext}}_{{\text{qm}}}({{q}}_{j}),\qquad\hat{\bf w}_{{\text{qm}}}=\sum_{j=1}^{N_{{\text{qm}}}}\sum_{i>j}{w}_{{\text{qm}}}({{q}}_{i},{{q}}_{j}).$ (4)
Then, ${\bf H}_{{\text{mm}}}$ is the classical Hamiltonian function representing $N_{{\text{mm}}}$ identical classical particles of mass $M$, in interaction with a local external potential $V^{\text{ext}}_{{\text{mm}}}({{Q}})$ and interacting among themselves with a local pairwise potential ${W}_{{\text{mm}}}({{Q}}_{i},{{Q}}_{j})$. These particles are typically effective particles such as atoms or rigid molecules instead of the individual electrons composing them and they interact through model, parametrized, force fields. The explicit form of the classical Hamiltonian ${\bf H}_{{\text{mm}}}$ reads then
${\bf H}_{{\text{mm}}}={\bf T}+{\bf V}^{\text{ext}}_{{\text{mm}}}+{\bf W}_{{\text{mm}}},$ (5)
with
${\bf T}=\sum_{i=1}^{N_{{\text{mm}}}}\frac{{{P}}_{i}^{2}}{2M},\qquad{\bf V}^{\text{ext}}_{{\text{mm}}}=\sum_{i=1}^{N_{{\text{mm}}}}V^{\text{ext}}_{{\text{mm}}}({{Q}}_{i}),$ (6)
${\bf W}_{{\text{mm}}}=\sum_{j=1}^{N_{{\text{mm}}}}\sum_{i>j}{W}_{{\text{mm}}}({{Q}}_{i},{{Q}}_{j}).$ (7)
Regarding the coupling term ${W}_{{\text{qm}}}^{{\text{mm}}}$, we simply assume here that it is an additive pairwise potential, i.e.
$\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}=\sum_{i=1}^{N_{{\text{qm}}}}\sum_{j=1}^{N_{{\text{mm}}}}{W}_{{\text{qm}}}^{{\text{mm}}}({{q}}_{i},{{Q}}_{j})$ (8)
The function ${W}_{{\text{qm}}}^{{\text{mm}}}({q},{Q})$ represents the effective interaction between an effective particle (such as a nuclei, an atom or a molecule) at position ${Q}$ and a fundamental particle (such as an electron) at position ${q}$.
II.2 Mixed quantum-classical equilibrium density
The equilibrium density corresponding to the Hamiltonian of Eq. (1) can be written as
$\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})=\frac{{\langle{\bf{q}}|}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}{|{\bf{q}}^{\prime}\rangle}e^{-\beta{\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})}}{Z^{\text{tot}}},$ (9)
where $\beta=(\text{k}_{\text{B}}T)^{-1}$, $\text{k}_{\text{B}}$ is the Boltzmann constant, and the dependency of $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}$ on the classical variables ${\bf{Q}}$ is here explicitly highlighted. This result is nothing but the real-space representation of that obtained by Nielsen et. al.Nielsen, Kapral, and Ciccotti (2001) in the adiabatic basis. The result of Eq. (9) can be obtained by starting from an Hamiltonian where both sets of particles are considered to be quantum. Then, one performs a Wigner transform over the $N_{{\text{mm}}}$ particles only and eventually a semi-classical limit is obtained by letting the ratio of masses $(m/M)^{1/2}$ tends towards 0. Conserving only the zeroth order term of this semi-classical limit yields to Eq. (9). This expansion in terms of the ratio of the masses is natural since the quantum particles, typically electrons, are usually much lighter than the classical ones, typically atoms or molecules. The original derivation of Nielsen et. al.Nielsen, Kapral, and Ciccotti (2001) used the Liouvillian imaginary time evolution equation of the density matrix together with the Poisson-brakets operatorsİmre et al. (1967), which is a well-known formalism in the quantum dynamics communityKapral and Ciccotti (1999); Nielsen, Kapral, and Ciccotti (2001); Kapral (2006, 2015). However, we provide in the supplementary materials an alternative derivation which involves a similarity transformation and resorts to the Baker-Campbell-Hausdorf expansion to express the Wigner transform Hamiltonian as a series of nested commutators, which might appear more familiar to the ”static” quantum chemistry community.
In Eq. (9), the normalization is the partition function of the QM/MM system which is expressed as
$Z^{\text{tot}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{N_{{\text{mm}}}!h^{3N_{{\text{mm}}}}}}\int\text{d}{\bf{Q}}\text{d}{\bf P}\!\!\int\!\text{d}{\bf{q}}$ ${\langle{\bf{q}}|}{e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}}{|{\bf{q}}\rangle}$ (10)
$e^{-\beta{\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})}.$
The corresponding Helmholtz free energy is given by
$F_{0}^{\text{tot}}=-\text{k}_{\text{B}}T\log(Z^{\text{tot}}).$ (11)
While the expressions of Eqs. (9) and (10) are given in real-space, they can easily be written in the adiabatic basis, where one explicitly sees that the density matrix is diagonal as found by Nielsen et. al.Nielsen, Kapral, and Ciccotti (2001).
Although the expression of the equilibrium density in Eq. (9) is relatively simple, it calls some comments. Firstly, as usual in classical statistical mechanics, the contribution to the Helmholtz free energy coming from the purely classical kinetic term in ${\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})$ can be computed analytically. Therefore it is worth splitting the equilibrium density in terms of ${\bf{Q}}$ and ${\bf P}$ dependence
$\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})=\frac{e^{-\beta{\bf T}}}{z_{{\bf P}}}\mathcal{P}_{\text{eq}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime}),$ (12)
where ${\bf T}$ is the purely classical kinetic energy defined in Eq. (6), and
$\mathcal{P}_{\text{eq}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})=\frac{{\langle{\bf{q}}|}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}{|{\bf{q}}^{\prime}\rangle}e^{-\beta{{\bf{V}}}_{\text{mm}}({\bf{Q}})}}{Z},$ (13)
${{\bf{V}}}_{\text{mm}}$ $={\bf V}^{\text{ext}}_{{\text{mm}}}+{\bf W}_{{\text{mm}}},$ (14)
$z_{{\bf P}}$ $=\frac{1}{N_{{\text{mm}}}!h^{3N_{{\text{mm}}}}}\int\text{d}{\bf P}e^{-\beta{\bf T}}=\frac{1}{N_{{\text{mm}}}!}\bigg{(}\frac{2\pi M}{\beta h^{2}}\bigg{)}^{\frac{3}{2}N_{{\text{mm}}}},$ (15)
$Z$ $=\int\text{d}{\bf{Q}}\int\text{d}{\bf{q}}{\langle{\bf{q}}|}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}{|{\bf{q}}\rangle}e^{-\beta{{\bf{V}}}_{\text{mm}}({\bf{Q}})},$ (16)
$Z^{\text{tot}}$ $=z_{{\bf P}}Z,$ (17)
such that the free energy can be written as
$F_{0}^{\text{tot}}$ $=F_{0}+F_{{\bf P}},$ (18)
where
$F_{{\bf P}}=-\text{k}_{\text{B}}T\log(z_{{\bf P}}),$ (19)
and
$F_{0}=-\text{k}_{\text{B}}T\log(Z)$ (20)
is the interacting part of the Helmholtz free energy of the QM/MM system.
Secondly, the object $\mathcal{P}_{\text{eq}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})$ in Eq.(13) can be expressed as the product of a purely quantum density matrix ${\boldsymbol{\rho}}_{\text{eq}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})$ (more precisely its matrix element evaluated in real space) and a purely classical function $\boldsymbol{p}_{\text{eq}}({\bf{Q}})$, i.e.
$\mathcal{P}_{\text{eq}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})={\boldsymbol{\rho}}_{\text{eq}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})\boldsymbol{p}_{\text{eq}}({\bf{Q}}),$ (21)
with
${\boldsymbol{\rho}}_{\text{eq}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})$ $={\langle{\bf{q}}|}\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}}){|{\bf{q}}^{\prime}\rangle},$ (22)
$\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}})=\frac{1}{Z^{{\text{qm}}}({\bf{Q}})}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))},$ (23)
$Z^{{\text{qm}}}({\bf{Q}})$ $=\int\text{d}{\bf{q}}{\langle{\bf{q}}|}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}{|{\bf{q}}\rangle},$ (24)
$\boldsymbol{p}_{\text{eq}}({\bf{Q}})=\frac{Z^{{\text{qm}}}({\bf{Q}})}{Z}e^{-\beta{{\bf{V}}}_{\text{mm}}({\bf{Q}})}.$ (25)
$Z$ and $Z^{{\text{qm}}}({\bf{Q}})$ are related through the following equation
$Z$ $=\int\text{d}{\bf{Q}}Z^{{\text{qm}}}({\bf{Q}})e^{-\beta{{\bf{V}}}_{\text{mm}}({\bf{Q}})}.$ (26)
It is noteworthy that $\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}})$ defined in Eq. (22),(23) is a positive definite operator and that $\boldsymbol{p}_{\text{eq}}({\bf{Q}})$ of Eq. (25) is also positive definite and normalized to unity. Therefore, $\hat{{\boldsymbol{\rho}}}_{\text{eq}}$ and $\boldsymbol{p}_{\text{eq}}$ are simply a special case of more general density matrices representing $N_{{\text{qm}}}$ quantum particles and probability distribution functions representing $N_{{\text{mm}}}$ classical particles, respectively. This will allow us in Sec. II.3 to propose a variational formulation of the Helmholtz free energy for a QM/MM system.
Additionally, this decomposition of the equilibrium density as a product of a purely quantum equilibrium density and a purely classical equilibrium density could lead to the somehow strange conclusion that there is no correlation between the quantum and classical systems, even if the two systems interact. Nevertheless, one must pay attention to the fact that the quantum part of the equilibrium density, $\hat{{\boldsymbol{\rho}}}_{\text{eq}}$, depends on the coordinates of the classical systems through the QM/MM interaction potential $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$. Therefore, the two systems are indeed correlated.
Thirdly, an obvious illustration of the meaning of the equilibrium density of Eq. (13) is that corresponding to the Born-Oppenheimer approximation. In the latter case, the light particles described by ${\bf{q}}$ are electrons and the heavy particles described by ${\bf{Q}}$ are the classical nuclei. There, $\hat{\bf h}_{{\text{qm}}}$ is composed of the kinetic operator of the electrons and their mutual coulomb interaction, there is no external potential both for the QM and MM parts, $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$ is the nuclei-electron attraction, and ${\bf W}_{{\text{mm}}}({\bf{Q}})$ is the Coulomb repulsion between the nuclei, such that the operator happening in the exponential of Eq. (13) is the Born-Oppenheimer Hamiltonian.
A follow up on this comment is that, as mentioned earlier, $\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}})$ is diagonal in the adiabatic basis, which is equivalent to say that non-adiabatic couplings are completely neglected. In the context of ab initio MD, this implies that $\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}})$ can be obtained by averaging, over all possible adiabatic electronic states, the classical trajectories of nuclei computed from the forces of a given adiabatic electronic state. Consequently, it entirely ignores surface hopping between different adiabatic states, which is why $\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}})$ is referred to as the adiabatic QM/MM density matrix. There is a flourishing literature concerned with going beyond this adiabatic approximation, specially in the context of quantum dynamics or statistical mechanics for QM/MM treatments (see for instance Ref. Kapral, 2015 for a review), but the treatment of non adiabatic couplings goes beyond our QM/MM scheme.
II.3 Reformulation of the Helmholtz free energy for a QM/MM system within a variational framework
II.3.1 Variational formulation
As we know that the equilibrium density for a mixed QM/MM system is expressed as a special case of product of a quantum density matrix by a classical probability distribution, we want to reformulate the Helmholtz free energy $F_{0}$ as a variational problem over QM and MM objects, such that the equilibrium density of Eq. (21) is the minimizer of that functional and that it yields the correct Helmholtz free energy of Eq. (20). Moreover, since the ultimate goal is to formulate a variational principle involving both quantum and classical densities, it is natural to use the decomposition of Eq. (21).
Thus, we would like to write the Helmholtz free energy as the minimum of a functional defined on a couple of quantum and classical objects as follows
$F_{0}=\min_{\hat{{\boldsymbol{\rho}}},\boldsymbol{p}}F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}],$ (27)
where the operator $\hat{{\boldsymbol{\rho}}}$ is a quantum density matrix representing $N_{{\text{qm}}}$ particles but also depending on all classical variables ${\bf{Q}}=\{{{Q}}_{1},\ldots,{{Q}}_{N_{{\text{mm}}}}\}$, i.e.
$\hat{{\boldsymbol{\rho}}}\equiv\hat{{\boldsymbol{\rho}}}({\bf{Q}}),$ (28)
and $\boldsymbol{p}$ is a classical probability density representing $N_{{\text{mm}}}$ classical particles and not depending on any quantum variable. Therefore, ${\boldsymbol{\rho}}$ and $\boldsymbol{p}$ must be normalized to be density matrices and probability measures, respectively, i.e
$\int\text{d}{\bf{q}}\,\,{\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})=1\quad\forall\,\,{\bf{Q}},$ $\int\text{d}{\bf{Q}}\,\,\boldsymbol{p}({\bf{Q}})=1,$ (29)
where ${\boldsymbol{\rho}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})$ is the real-space representation of the operator $\hat{{\boldsymbol{\rho}}}$, i.e.
${\boldsymbol{\rho}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})={\langle{\bf{q}}|}\hat{{\boldsymbol{\rho}}}({\bf{Q}}){|{\bf{q}}^{\prime}\rangle},$ (30)
such that that the product of the two functions is normalized to unity
$\text{Tr}\{\hat{{\boldsymbol{\rho}}}\boldsymbol{p}\}=\int\text{d}{\bf{Q}}\int\text{d}{\bf{q}}\,\,{\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})\boldsymbol{p}({\bf{Q}})=1.$ (31)
Therefore, the equilibrium density of the mixed QM/MM system described in Eq. (21) lies within the functional space defined by all the couple of functions and density matrices satisfying Eqs. (29)-(31). As shown in the appendix (see section Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$), if we define the Helmholtz functional as follows
$F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=\mathcal{U}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]-\text{k}_{\text{B}}T\mathcal{S}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}],$ (32)
where $\mathcal{U}$ is an energy functional excluding the classical kinetic term (which has was accounted for in $F_{{\bf P}}$)
$\mathcal{U}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=\text{Tr}\big{\{}\hat{{\boldsymbol{\rho}}}\boldsymbol{p}(\hat{\bf h}_{{\text{qm}}}+\hat{{\bf v}}^{\text{ext}}_{{\text{qm}}}+{\bf W}_{{\text{mm}}}+{\bf V}^{\text{ext}}_{{\text{mm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}})\big{\}},$ (33)
and $\mathcal{S}$ is the dimensionless entropy of the total system
$\mathcal{S}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=-\text{Tr}\big{\{}\hat{{\boldsymbol{\rho}}}\boldsymbol{p}\log\big{(}\hat{{\boldsymbol{\rho}}}\boldsymbol{p}\big{)}\big{\}},$ (34)
then the only minimizer of the functional of Eq. (32) is $(\hat{{\boldsymbol{\rho}}}_{\text{eq}},\boldsymbol{p}_{\text{eq}})$ of Eq. (21), i.e
$-\text{k}_{\text{B}}T\log(Z)\leq F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]\quad\forall\,\,(\hat{{\boldsymbol{\rho}}},\boldsymbol{p}),$ (35)
and
$F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=F_{0}=-\text{k}_{\text{B}}T\log(Z)\Leftrightarrow(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})=(\hat{{\boldsymbol{\rho}}}_{\text{eq}},\boldsymbol{p}_{\text{eq}}).$ (36)
II.3.2 Rewriting of the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$
Because $\boldsymbol{p}$ does not depend on the quantum variables and that we chose the normalization condition of Eq. (29), we can naturally decompose the entropy as
$\mathcal{S}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ $=S_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]+S_{{\text{mm}}}[\boldsymbol{p}],$ (37)
with $S_{{\text{mm}}}[\boldsymbol{p}]$ being the purely classical dimension-less entropy
$S_{{\text{mm}}}[\boldsymbol{p}]$ $=-\int\text{d}{\bf{Q}}\,\boldsymbol{p}({\bf{Q}})\log(\boldsymbol{p}({\bf{Q}})),$ (38)
and $S_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ is the quantum-classical dimension-less entropy
$S_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ $=-\text{Tr}\big{\{}\hat{{\boldsymbol{\rho}}}\boldsymbol{p}\log(\hat{{\boldsymbol{\rho}}})\big{\}}$ (39)
$=-\int\text{d}{\bf{Q}}\,\boldsymbol{p}({\bf{Q}})\int\text{d}{\bf{q}}\,{\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})\log({\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})),$
which depends explicitly on both $\hat{{\boldsymbol{\rho}}}$ and $\boldsymbol{p}$.
We can similarly decompose $\mathcal{U}$ as follows
$\mathcal{U}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ $=E_{{\text{mm}}}[\boldsymbol{p}]+E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ (40)
$+E_{\text{qm}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]+E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]+E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}],$
where $E_{\text{mm}}[\boldsymbol{p}]$ is the mutual interaction of the classical part
$E_{\text{mm}}[\boldsymbol{p}]$ $=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}}){\bf W}_{{\text{mm}}}({\bf{Q}}),$ (41)
$E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ is the functional accounting for the interaction of the classical particles with the external potential
$E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ $=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}}){\bf V}^{\text{ext}}_{{\text{mm}}}({\bf{Q}}),$ (42)
$E_{\text{qm}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ is the kinetic and mutual interaction of the quantum part
$E_{\text{qm}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ $=\text{Tr}\{(\hat{\bf t}+\hat{\bf w}_{{\text{qm}}})\boldsymbol{p}\hat{{\boldsymbol{\rho}}}\}$ (43)
$=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}})$
$\quad\int\text{d}{\bf{q}}\text{d}{\bf{q}}^{\prime}(\bf{t}({\bf{q}},{\bf{q}}^{\prime})+{w}_{{\text{qm}}}({\bf{q}},{\bf{q}}^{\prime})){\boldsymbol{\rho}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}}),$
$E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ is the functional accounting for the interaction of the quantum particles with the external potential
$E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ $=\text{Tr}\{\hat{{\bf v}}^{\text{ext}}_{{\text{qm}}}\boldsymbol{p}\hat{{\boldsymbol{\rho}}}\}$ (44)
$=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}})\int\text{d}{\bf{q}}{\bf v}^{\text{ext}}_{{\text{qm}}}({\bf{q}},{\bf{q}}){\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}}),$
and eventually $E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ is the quantum-classical interaction,
$E_{{\text{qm}}}^{{\text{mm}}}$ $[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=\text{Tr}\{\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}\boldsymbol{p}\hat{{\boldsymbol{\rho}}}\}$ (45)
$=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}})\int\text{d}{\bf{q}}\int\text{d}{\bf{q}}^{\prime}{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime}){\boldsymbol{\rho}}({\bf{q}}^{\prime},{\bf{q}},{\bf{Q}}).$
It is noteworthy here to highlight that the form of Eq. (45) is quite general as it consists in a priori in potential with a non local kernel ${\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})$. This form of interaction includes for instance Lennard-Jones potentials, frozen density Coulomb interaction or density overlap models for Pauli repulsionWheatley and Price (1990).
Eventually, once all terms are inserted in Eq. (27), the Helmholtz free energy $F_{0}$ can be obtained as
[…]
$+E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]-\text{k}_{\text{B}}T\big{(}S_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]+S_{{\text{mm}}}[\boldsymbol{p}]\big{)}\big{\}}.$
Eq. (46) is therefore a variational formulation of $F_{0}$ for a mixed QM/MM system in terms of a minimization over the couple $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})$, which, to the best of our knowledge, was never formalized before. One can see from Eqs. (38), (41) and (42) that the functionals corresponding to the MM part of the energy and entropy are completely independent of the quantum part of the system, as they depend only on the classical probability $\boldsymbol{p}$. Nevertheless, because of the term $E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$, the quantum and classical parts are coupled in the minimization process of Eq. (46). Also, one can see that the functionals corresponding to the QM parts of the energy and entropy explicitly depend on the classical MM probability distribution $\boldsymbol{p}$.
Although Eq. (46) provides a variational formulation of the Helmholtz free energy for a QM/MM system, the variables $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})$ involved in the minimization are nevertheless very high-dimensional. Therefore, the goal of the following sections is to reformulate the variational principle of Eq. (46) using functionals of significantly lower-dimensional densities, i.e. the one-body QM and MM densities.
III One-body density formulation through Levy-Lieb constrained search formulation
An important consequence of the one-body nature of both $V^{\text{ext}}_{{\text{mm}}}({Q})$ and $v^{\text{ext}}_{{\text{qm}}}({q})$, along with the permutational symmetry of both $\boldsymbol{p}$ and $\hat{{\boldsymbol{\rho}}}$ is that the corresponding energy terms $E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ and $E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ in Eqs. (42) and (44) can be computed using one-body objects. Therefore, by applying the Levy-Lieb constrained search, we can reformulate the variational problem of Eq. (46) in terms of a functional that depends only on the one-body quantum and classical densities, rather than the full $N_{{\text{mm}}}+N_{{\text{qm}}}$ variables.
III.1 Classical and quantum one-body density and the Levy-Lieb constrained search
We naturally introduce the one-body classical density $n({Q})$ as the trace over all quantum variables and all classical variables but one of the product $\hat{{\boldsymbol{\rho}}}({\bf{q}},{\bf{Q}})\boldsymbol{p}({\bf{Q}})$, which, thanks to the normalization of Eq. (29), is simply
[…]
Similarly we can define the QM one-body density $\rho({q})$ as
[…]
It is noteworthy from Eq. (48) that, because of the ${\bf{Q}}$ dependence in $\hat{{\boldsymbol{\rho}}}({\bf{Q}})$, the classical $N$-body probability $\boldsymbol{p}({\bf{Q}})$ contributes to the one-body quantum density $\rho({q})$, whereas from Eq. (47) it is clear that the quantum $N$-body density matrix $\hat{{\boldsymbol{\rho}}}({\bf{Q}})$ does not contribute to the classical one-body density.
By using permutation symmetry of both $\hat{{\boldsymbol{\rho}}}({\bf{Q}})$ and $\boldsymbol{p}({\bf{Q}})$ with respect to exchange of either QM or MM variables, we can then write the two terms arising from the external potential as functionals of the QM and MM one-body densities
$E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ $=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}}){\bf V}^{\text{ext}}_{{\text{mm}}}({\bf{Q}})$ (49)
$=\int\text{d}{Q}n({Q})V^{\text{ext}}_{{\text{mm}}}({Q})$
$\equiv(V^{\text{ext}}_{{\text{mm}}}|n),$
$E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ $=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}})\int\text{d}{\bf{q}}{\bf v}^{\text{ext}}_{{\text{qm}}}({\bf{q}},{\bf{q}}){\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})$ (50)
$=\int\text{d}{q}\rho({q})v^{\text{ext}}_{{\text{qm}}}({q})$
$\equiv(v^{\text{ext}}_{{\text{qm}}}|\rho).$
As the contributions coming from the external potentials can be computed directly with a couple of one-body object $(\rho,n)$, we will now introduce a universal functional of $(\rho,n)$ which contains all necessary information to compute the Gibbs free energy. This functional is built using the Levy-Lieb constrained search formalism as in eDFT. First, we introduce the intrinsic Gibbs free energy functional of the $N$-body objects
$f[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ $=E_{{\text{mm}}}[\boldsymbol{p}]+E_{\text{qm}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]+E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]-\text{k}_{\text{B}}T\mathcal{S}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}].$ (51)
such that we can then rewrite the Gibbs free energy as
$F_{0}=\min_{\hat{{\boldsymbol{\rho}}},\boldsymbol{p}}\big{(}E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]+E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]+f[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]\big{)}.$ (52)
Then, the many-to-one correspondence between $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})$ and $(\rho,n)$ allows us to introduce the Levy-Lieb constrained search as follows
$\min_{\hat{{\boldsymbol{\rho}}},\boldsymbol{p}}F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=\min_{(\rho,n)}\,\,\min_{(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})\rightarrow(\rho,n)}F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}],$ (53)
where the notation $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})\rightarrow(\rho,n)$ means that the search is done over the set of couple of $N$-body objects $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})$ having the couple of 1-body object given by $(\rho,n)$. Therefore, by remembering that the external parts depends only on the one-body densities (see Eqs. (49) and (50)), one can rewrite Eq. (53) as
$\min_{\hat{{\boldsymbol{\rho}}},\boldsymbol{p}}F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=$ $\min_{(\rho,n)}\big{[}(V^{\text{ext}}_{{\text{mm}}}|n)+(v^{\text{ext}}_{{\text{qm}}}|\rho)$ (54)
$+$ $\min_{(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})\rightarrow(\rho,n)}f[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]\big{]}.$
Therefore, we can introduce the universal Levy-Lieb functional of the one-body classical and quantum density as
$\mathcal{F}[\rho,n]=\min_{(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})\rightarrow(\rho,n)}f[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}],$ (55)
such that one can rewrite the Gibbs free energy as
$F_{0}=\min_{(\rho,n)}\big{(}(V^{\text{ext}}_{{\text{mm}}}|n)+(v^{\text{ext}}_{{\text{qm}}}|\rho)+\mathcal{F}[\rho,n]\big{)},$ (56)
which achieves to prove that we can obtain $F_{0}$ as a minimization over only one-body densities.
III.2 Explicit coupling in terms of eDFT and cDFT
Although Eq. (56) allows us to obtain the Gibbs free energy only in terms of the one-body densities $(\rho,n)$, such a functional remains unknown and the explicit connection with well known universal quantum and classical density functionals is yet unclear. This paragraph is concerned with a rewriting of $\mathcal{F}[\rho,n]$ such that the functional of MerminMermin (1965) and EvansEvans (1979) explicitly appear, together with a new QM/MM correlation functional.
We begin by recalling the universal purely quantum one-body density functionals of Mermin
$\mathcal{F}_{\text{qm}}[\rho]=\min_{\hat{{\boldsymbol{\rho}}}\rightarrow\rho}\text{Tr}\{\hat{{\boldsymbol{\rho}}}\big{(}\hat{\bf t}+\hat{\bf w}_{{\text{qm}}}+\text{k}_{\text{B}}T\log(\hat{{\boldsymbol{\rho}}})\big{)}\},$ (57)
where $\hat{{\boldsymbol{\rho}}}$ is a quantum $N$-body density matrix and $(\hat{{\boldsymbol{\rho}}}\rightarrow\rho)$ means that the minimization is performed over $N$-body quantum density matrices $\hat{{\boldsymbol{\rho}}}$ leading to a given quantum one-body density $\rho$. Similarly, we can define the universal classical one-body functional of Evans
$\mathcal{F}_{\text{mm}}[n]=\min_{\boldsymbol{p}\rightarrow n}\text{Tr}\{\boldsymbol{p}\big{(}{\bf W}_{{\text{mm}}}+\text{k}_{\text{B}}T\log(\boldsymbol{p})\big{)}\},$ (58)
where $\boldsymbol{p}$ is a classical $N$-body probability distribution and $(\boldsymbol{p}\rightarrow n)$ means that the minimization is performed over $N$-body classical probability distribution $\boldsymbol{p}$ leading to a given classical one-body density $n$. Historically, both Mermin and Evans introduced their universal functionals in the Grand Canonical ensemble, so Eqs. (57) and (58) are actually the Canonical counter part of the original functionals. Also, in their respective original papersMermin (1965); Evans (1979), neither Mermin or Evans explicitly used the constrained search framework which was formalized later by Levy and LiebLevy (1979); Lieb (1983).
We can then rewrite $\mathcal{F}[\rho,n]$ as follows
$\mathcal{F}[\rho,n]=$ $\mathcal{F}_{\text{qm}}[\rho]+\mathcal{F}_{\text{mm}}[n]+\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]+\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n],$ (59)
where $\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ is the mean-field QM/MM interaction
$\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ $=\int\text{d}{q}\int\text{d}{Q}{W}_{{\text{qm}}}^{{\text{mm}}}({Q},{q})n({Q})\rho({q}),$ (60)
which will be the QM/MM equivalent of the Hartree-term in eDFT, and where $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ is the QM/MM correlation functional defined of course as the complementary functional
$\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]=\mathcal{F}[\rho,n]-\big{(}\mathcal{F}_{\text{qm}}[\rho]+\mathcal{F}_{\text{mm}}[n]+\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]\big{)}.$ (61)
The functional $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ in Eq. (61) is there to reproduce all correlation terms arising because of the QM/MM interaction. The latter contains therefore some energetic and entropic terms.
We can then rewrite explicitly $F_{0}$ as
[…]
$+\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]+\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]\big{\}},$
where we can explicitly recognize the purely quantum and purely classical free-energy functionals, together with the QM/MM coupling expressed as the sum of the mean-field and correlation terms. If $\delta\mathcal{F}_{{\text{qm}}}$ is ignored, we recover the mean-field coupling between the eDFT and cDFT functionals that is generally used in the literaturePetrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007); Tang et al. (2020); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024).
III.3 Applications to solvation problems
The development presented here were motivated by the necessity of a well grounded theoretical framework for a DFT description of a quantum solute solvated in a classical solvent. This section discusses the application of Eq. (62) to this particular problem.
As common in electronic DFT calculations with an implicit solvent description, the nuclei of the molecular solute are considered clamped. The Hamiltonian of the quantum system of Eq. (2) is therefore the Born-Oppenheimer Hamiltonian.
Regarding now the classical Hamiltonian, it strongly depends on the nature of the considered solvent. We simply assume here that the external part of this potential, ${\bf V}^{\text{ext}}_{{\text{mm}}}$, represents the interaction between the solvent and the classical nuclei of the solute, and that it is local and expressed as a sum of one-body terms. Therefore, ${\bf V}^{\text{ext}}_{{\text{mm}}}$ naturally contains a Coulombic term but additional terms can be incorporated, in order, for instance, to model repulsion and dispersion interactions. There exists multiple types of classical potential modeling dispersion and repulsion, the most popular one being the Lennard-Jones potential.
Central to the description of the QM/MM interaction, $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}$, accounts for the electron-solvent interaction. This is a part of the total Hamiltonian that is both unknown and whose quality is critical since it essentially accounts for the information that has been lost when replacing quantum particles by classical particles. This interaction is often modeled using a Coulomb potential. Since only the external potentials are required to be local to derive Eq. (62), more sophisticated models incorporating non-local quantum operators can also be incorporated in $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}$. This is a route that should be explored to improve the quality of the QM/MM model.
We now assume that, in the absence of the solvent, the typical electronic gap is large compared to the thermal agitation $\text{k}_{\text{B}}T$, allowing us to neglect the contribution of electronic entropy in $\mathcal{F}_{\text{qm}}[\rho]$ (see Eq. (57)). This is equivalent to considering only the ground state of the electronic quantum system. To illustrate the validity of this approximation for systems at room temperature ($T=300\,K$), we note that the Boltzmann weight ratio between the ground state and an excited state with a gap of 0.25 eV ($\approx 0.01\,\text{Ha}$) is approximately $10^{-5}$. For a gap of 1.63 eV ($\approx 0.06\,\text{Ha}$), which marks the onset of the visible spectrum, this ratio drops dramatically to $10^{-28}$. Since most chemical systems near their equilibrium geometry exhibit electronic gaps of several eV, this approximation is well-suited for typical room-temperature conditions. As a consequence, we can take the limit $T\rightarrow 0$ in the definition of the Mermin universal functional (see Eq. (57)) to recover the usual universal Levy-Lieb electronic density functional, namely
$\mathcal{F}_{\text{qm}}[\rho]\approx\min_{\Psi\rightarrow\rho}{\langle\Psi|}\hat{\bf t}+\hat{\bf w}_{{\text{qm}}}{|\Psi\rangle}.$ (63)
Regarding solvent effects, we focus on cases where the solvent-solute interaction is not too strong, meaning there is no covalent bonding between the solute and solvent species. Under these conditions, the previous $T=0~{}K$ assumption for $\mathcal{F}_{\text{qm}}[\rho]$ remains reasonable, as the electronic structure of the solute is not significantly altered by the presence of the solvent.
In previous work about mixed electronic-classical density functional theory to study QM/MM systems, the classical fluid was described in the Grand Canonical ensemble. This is because classical DFT is naturally formulated in this ensemble. In practice this means that the number of classical particle is now allowed to fluctuate by exchanging with a reservoir of fixed chemical potential $\mu$. In this ensemble, Eq. (62)
[…]
$+\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]+\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]\big{\}}$
where $\Omega_{0}$ is the grand potential, that is the state function in the grand canonical ensemble, and $\mathcal{F}_{\text{mm}}$ is now the grand canonical universal functional a la Evans.
In Eq. (64) one recognizes the usual quantum Born-Oppenheimer part for the electron (i.e. $(v_{\text{ne}}|\rho)+\mathcal{F}_{\text{qm}}[\rho]$), together with the purely classical part of the solvent (i.e. $\mathcal{F}_{\text{mm}}[n]+(v_{nucl}^{{\text{mm}}}-\mu|n)$), and $\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ accounts for the mean-field electron-solvent interaction, while $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ accounts for solvation correlation effects in the QMMM system.
In practice, Eq. (64) requires some approximations. First, the choice of the functionals describing the quantum system and the liquid considerably influences the quality of the predictions. However, determining the appropriate electronic functional to describe a given solute and the suitable functional for a classical fluid are topics that themselves deserve a thorough review. Therefore, we let this discussion aside to focus on the QM/MM coupling. In Petrosyan et al’s joint DFT Petrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007), the interaction of the solvent with both the solute nuclei and electrons is dealt as a whole and the coupling is described by the use of pseudo potential. Thus Eq. (64) solely contains the quantum functional $\mathcal{F}_{\text{qm}}$, the solvent one $\mathcal{F}_{\text{mm}}$ and $\varepsilon_{{\text{qm}}}^{{\text{mm}}}$. The coupling term is modeled as a dielectric screening due to a non local dielectric function depending on the electron density and a mean-field term between both densities interacting through a Gaussian repulsive potential.
In Tang et alTang et al. (2020) Reaction DFT, both electronic and classical densities are not optimized simultaneously. Instead, the quantum functional is optimized in a continuum model. The ground state electronic density is used to compute point charges on the solute atoms which generate an external electrostatic potential in the subsequent classical functional optimization. The interaction between the quantum and the classical parts is similar to Petrosyan’s work, i.e the contribution of the electron and nuclei are gathered together. It is limited to the mean-field term, here a Coulombic term between point charges and a Lennard-Jones term and the correlation functional $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}$ is lacking. Moreover, since both density are not optimized self-consistently the interaction term $\varepsilon_{{\text{qm}}}^{{\text{mm}}}$ only influences the classical optimization and not the quantum one.
In our previous worksJeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024), we also ignored the correlation functional $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}$ in Eq. (62) and limit ourselves to the mean-field term. The interaction potential was also the sum of a Lennard-Jones and a Coulombic term. However, we do not resort to point charges representation to compute the electrostatic energy but directly use the electronic density
[…]
where $n_{C}$ is the solvent charge density.
We also proposed to couple wave function theory-based methods, such as Hartree-Fock and selected configuration interaction with cDFTLabat, Giner, and Jeanmairet (2024) to approximate the universal Levy-Lieb functional $\mathcal{F}_{\text{qm}}[\rho]$ of Eq.62.
The results obtained in these worksPetrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007); Tang et al. (2020); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024) clearly indicate that, for ground state properties such as dipoles for instance, the mean-field QM/MM approximation gives already a decent improvement with respect to in vaccuum pure QM calculations.
IV Conclusion
In the present work we have developed exact variational formulations of the Helmholtz free energy of a QM/MM system made of $N_{{\text{qm}}}$ and $N_{{\text{mm}}}$ identical particles. An important result is that the exact Helmholtz free energy of a QM/MM system can be obtained though a minimization of functionals over objects of reduced dimensions, namely the one-body classical and quantum density.
The starting point of these derivations is the proposal of the QM/MM equivalent to the quantum ground state variational principle or the classical Gibbs measure variational principle (see Sec. II.3). This QM/MM functional is formalized in terms of a $N_{{\text{mm}}}$ classical probability density and an $N_{{\text{qm}}}$ quantum density matrix which also parametrically depends on the $N_{{\text{mm}}}$ classical variables (see Sec. Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ for proofs of the variational property). The latter was proposed thanks to the knowledge of the equilibrium density of the QM/MM system (see Eqs. (13) and (16)) which can be obtained from first principle through a semi classical expansion of a partial Wigner transformation (see Ref. Nielsen, Kapral, and Ciccotti, 2001 and the section II of the supplementary information for a detailed derivation).
Starting from this QM/MM variational principle written in terms of $N$-body objects and thanks to constrained search formalisms similar to that proposed by Levy and Lieb in eDFT, we establish a universal QM/MM functional depending only on the one-body classical and quantum objects (see Sec. III.1). Building on these results, we explicitly make a coupling with the finite temperature eDFT and cDFT functionals, which allows us to rigorously link the two theories for the description of a QM/MM system (see Sec. III.2). We then propose a practical application of this theory in the case of solvation problems, and propose a mean-field treatment. Thanks to these derivations, it becomes clear that the explicit correlation between the classical and quantum electronic systems, were implicitly neglected in our previously reported QM/cDFT calculationsJeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024).
Supplementary materials
The supplementary materials document contains additional information concerning: I) a recall of the variational formulation of statistical quantum mechanics in the canonical ensemble and II) an alternative derivation of the adiabatic QM-MM equilibrium density which does not rely on the Liouville equations or Poisson bracket operators.
Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$
We wish to show that i) the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ of Eq. (32) is bounded from below by the Helmholtz free energy $F_{0}=-\text{k}_{\text{B}}T\log(Z)$ with $Z$ defined in Eq. (16), and ii) that the only minimizer of $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ is $\mathcal{P}_{\text{eq}}$ of in Eq. (13). We will use Jensen’s inequality from convex analysis and probability theory to prove these two statements straightforwardly. This path is similar to that proposed by Friedli and Velenik in the case of classical lattice systems (see Lemma 6.74 in Ref. Friedli and Velenik, 2017) but here applied to the mixed QM/MM case. In order to be able to use Jensen’s inequality, we will first need to show that the couple $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})$ is indeed a probability measure. While for $\boldsymbol{p}$ it is obvious, it requires a little more attention for $\hat{{\boldsymbol{\rho}}}$ as it consists formally in a operator. Then, $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ will be written as an average value on probability measures, and Jensen’s inequality and strict convexity will do the rest.
.1 Rewriting $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ as an average value on probability measures
We use the adiabatic basis ${{\mathcal{B}}({\bf{Q}})=\{{|\Psi_{i}({\bf{Q}})\rangle}},0\leq i\leq\infty\}$ to develop the quantum or mixed quantum-classical operators. More precisely, ${\mathcal{B}}({\bf{Q}})$ consists in the complete set of eigenvectors of the following Schrodinger equation
$\big{(}\hat{{\bf h}}^{\text{tot}}_{\text{qm}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})\big{)}{|\Psi_{i}({\bf{Q}})\rangle}=E_{i}({\bf{Q}}){|\Psi_{i}({\bf{Q}})\rangle},$ (A.1)
such that the Hamiltonian $\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$ is naturally written as
$\hat{{\bf h}}^{\text{tot}}_{\text{qm}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})=\sum_{i}E_{i}({\bf{Q}}){|\Psi_{i}({\bf{Q}})\rangle}{\langle\Psi_{i}({\bf{Q}})|},$ (A.2)
and the operator $\hat{{\boldsymbol{\rho}}}({\bf{Q}})$ is written as follows
$\hat{{\boldsymbol{\rho}}}({\bf{Q}})=\sum_{i}{\boldsymbol{\rho}}_{ii}({\bf{Q}}){|\Psi_{i}({\bf{Q}})\rangle}{\langle\Psi_{i}({\bf{Q}})|}.$ (A.3)
Because of the properties of density matrices, the real-numbers ${\boldsymbol{\rho}}_{ii}({\bf{Q}})$ fulfill
${\boldsymbol{\rho}}_{ii}({\bf{Q}})\geq 0,\sum_{i\in{\mathcal{B}}({\bf{Q}})}{\boldsymbol{\rho}}_{ii}({\bf{Q}})=1.$ (A.4)
Therefore, as $\boldsymbol{p}({\bf{Q}})$ is a probability distribution and because of the properties of Eq. (A.4), if we define the function $\mu({\bf{Q}},i)$ as
${\mathbb{R}}^{3N_{{\text{qm}}}}\times{\mathbb{N}}\rightarrow{\mathbb{R}},$ (A.5)
$({\bf{Q}},i)\mapsto\mu({\bf{Q}},i)=\boldsymbol{p}({\bf{Q}}){\boldsymbol{\rho}}_{ii}({\bf{Q}}),$
then $\mu({\bf{Q}},i)$ fulfils the requirements for a probability measure, i.e.
$\mu({\bf{Q}},i)\geq 0\,\,\forall\,\,({\bf{Q}},i),\quad\int\text{d}{\bf{Q}}\sum_{i}\mu({\bf{Q}},i)=1.$ (A.6)
For the sake of compactness, we define ${\mathbf{X}}=({\bf{Q}},i)$, such that we can rewrite the probability measure as
$\mu({\bf{Q}},i)\equiv\mu({\mathbf{X}}),\quad\int\text{d}{\bf{Q}}\sum_{i}\equiv\int\text{d}{\mathbf{X}}.$ (A.7)
The equilibrium density $\hat{\mathcal{P}}_{\text{eq}}$ of Eq. (13) can be written using the adiabatic basis of Eq. (A.2) as follows
$\hat{\mathcal{P}_{\text{eq}}}({\bf{Q}})$ $=\frac{1}{Z}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}e^{-\beta\hat{{\bf V}}^{\text{tot}}_{\text{mm}}({\bf{Q}})}$ (A.8)
$=\frac{1}{Z}\sum_{i}e^{-\beta\epsilon({\bf{Q}},i)}{|\Psi_{i}({\bf{Q}})\rangle}{\langle\Psi_{i}({\bf{Q}})|},$
which corresponds to the measure $\mu({\mathbf{X}})=\mu_{\text{eq}}({\mathbf{X}})$ defined as follows
$\mu_{\text{eq}}({\mathbf{X}})\equiv\frac{e^{-\beta\epsilon({\mathbf{X}})}}{Z},$ (A.9)
where
$\epsilon({\mathbf{X}})=\epsilon({\bf{Q}},i)\equiv E_{i}({\bf{Q}})+\hat{{\bf V}}^{\text{tot}}_{\text{mm}}({\bf{Q}}).$ (A.10)
We introduce the notation for an average value over a probability measure $\mu({\mathbf{X}})$ as
$\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})f({\mathbf{X}})=\langle f\rangle_{\mu},$ (A.11)
such that the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ can be rewritten as an average value on the measure $\mu({\mathbf{X}})$
$F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=$ $\text{Tr}\big{\{}\hat{{\boldsymbol{\rho}}}\boldsymbol{p}\big{(}\hat{{\bf V}}^{\text{tot}}_{\text{mm}}+\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}+\text{k}_{\text{B}}T\log(\hat{{\boldsymbol{\rho}}}\boldsymbol{p})\big{)}\}$ (A.12)
$=$ $\int\text{d}{\bf{Q}}\sum_{i}\mu({\bf{Q}},i)\big{(}\epsilon({\bf{Q}},i)+\text{k}_{\text{B}}T\log(\mu({\bf{Q}},i))$
$=$ $\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\big{(}\epsilon({\mathbf{X}})+\text{k}_{\text{B}}T\log(\mu({\mathbf{X}}))$
$=$ $-\text{k}_{\text{B}}T\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\log(\frac{e^{-\beta\epsilon({\mathbf{X}})}}{\mu({\mathbf{X}})})$
$=$ $-\text{k}_{\text{B}}T\langle\log(\frac{e^{-\beta\epsilon}}{\mu})\rangle_{\mu}$
$\equiv$ $F[\mu].$
According to Eq. (A.12), the evaluation $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ at $\mathcal{P}_{\text{eq}}$ leads straightforwardly to
$F[\mu_{\text{eq}}]=$ $-\text{k}_{\text{B}}T\log(Z),$ (A.13)
which is of course the correct Helmholtz free energy. Then two questions remain: i) is $-\text{k}_{\text{B}}T\log(Z)$ a global minima and ii) is $\hat{\mathcal{P}_{\text{eq}}}$ the only minimizer ?
.2 Use of Jensen inequality to prove variational character
As the function $-\log(x)$ is strictly convex, we will now use Jensen’s inequality from convex analysis and probability measures to prove that $F[\mu]$ is bounded from below by the Helmholtz free energy and that $\mu_{\text{eq}}$ is the only minimizer. The Jensen’s inequality can be stated as follows.
If $\varphi$ is a convex function (i.e. always above its tangent lines), $\mu({\mathbf{X}})$ a probability measure on ${\mathbf{X}}\in E$ where $E$ is a normed vector space, $g({\mathbf{X}})$ a function from $E$ to ${\mathbb{R}}$, then Jensen’s equality states that
$\varphi(\langle g\rangle_{\mu})\leq\langle\varphi(g)\rangle_{\mu},$ (A.14)
or written in terms of integrals
$\varphi\bigg{(}\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})g({\mathbf{X}})\bigg{)}\leq\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\varphi\big{(}g({\mathbf{X}})\big{)}.$ (A.15)
As $-\log(x)$ is convex, the Jensen’s equality can be applied to $F[\mu]$ of Eq. (A.12) by setting
$g({\mathbf{X}})=\frac{e^{-\beta\epsilon({\mathbf{X}})}}{\mu({\mathbf{X}})},$ (A.16)
$\varphi(x)=-\text{k}_{\text{B}}T\log(x).$
Therefore the inequality of Eq. (A.14) becomes
$-\text{k}_{\text{B}}T\log\langle\frac{e^{-\beta\epsilon}}{\mu}\rangle_{\mu}\leq-\text{k}_{\text{B}}T\langle\log(\frac{e^{-\beta\epsilon}}{\mu})\rangle_{\mu},$ (A.17)
and as
$\langle\frac{e^{-\beta\epsilon}}{\mu}\rangle_{\mu}$ $=\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\frac{e^{-\beta\epsilon(x))}}{\mu({\mathbf{X}})}$ (A.18)
$=Z,$
one can rewrite Eq. (A.17) as
$-\text{k}_{\text{B}}T\log(Z)\leq F[\mu],$ (A.19)
which guarantees the variational character. Eventually, the strict convexity of the $-\log(x)$ function implies that the Jensen’s inequality becomes an equality if and only if $g({\mathbf{X}})=cst$, which translates into
$g({\mathbf{X}})=cst\Leftrightarrow\frac{e^{-\beta\epsilon({\mathbf{X}})}}{\mu({\mathbf{X}})}=cst\Leftrightarrow\mu({\mathbf{X}})\propto e^{-\beta\epsilon({\mathbf{X}})},$ (A.20)
therefore implying that $\mu_{\text{eq}}({\mathbf{X}})$ is the only minimizer. We point out that this proof essentially consists in the Gibb’s inequality re-framed with convex analysis vocabulary.