### v5

# A variational formulation of the free energy of mixed quantum-classical systems: coupling classical and electronic density functional theories

Abstract
Combining classical density functional theory (cDFT) with quantum mechanics (QM) methods offers a computationally efficient alternative to traditional QM/molecular mechanics (MM) approaches for modeling mixed quantum-classical systems at finite temperatures. However, both QM/MM and QM/cDFT rely on somewhat ambiguous approximations , the two major ones being: i) the definition of the QM and MM regions as well as the description of their coupling, and ii) the choice of the methods and levels of approximation made to describe each region
This paper addresses the second point and develop an exact theoretical framework that allows us to clarify the approximations involved in the QM/cDFT formulation. We therefore establish a comprehensive density functional theory (DFT) framework for mixed quantum-classical systems within the canonical ensemble. We start by recalling the expression of the adiabatic equilibrium density matrix for a mixed system made of $N_{{\text{qm}}}$ quantum and $N_{{\text{mm}}}$ classical particles. Then, we propose a variational formulation of the Helmholtz free energy in terms of the full , non-equilibrium, QM/MM density matrix. Taking advantage of permutational symmetry and thanks to constrained-search methods, we reformulate the computation of the Helmholtz free energy using only the quantum and classical one-body densities. Therefore, this paper generalizes both cDFT and electronic DFT (eDFT) to QM/MM systems.
We then reformulate the functional to make the standard eDFT and cDFT Levy–Lieb functionals explicitly appear, together with a new universal correlation functional for QM/MM systems. A mean-field approximation is finally introduced in the context of solvation problems and we discuss its connection with several existing mixed cDFT-eDFT schemes. An extension to the semi-grand canonical ensemble, where the number of classical particles is allowed to fluctuate, is provided in the supplementary materials.
I Introduction
The Hohenberg-Kohn theoremsHohenberg and Kohn (1964) dramatically reduce the complexity of simulating a quantum (QM) molecular system by enabling a focus on the one-electron density instead of the full $N$-electron wave function. Central to this simplification is the unknown universal density functional, which was later given a more rigorous mathematical foundation by Levy and LiebLevy (1979); Lieb (1983). Through the self-consistent framework introduced by Kohn and ShamKohn and Sham (1965), along with extensive efforts to develop approximations to the exact exchange-correlation functional (see, for example, a recent review in Ref. Toulouse, 2022), electronic density functional theory (eDFT) has become the workhorse of quantum chemistry.
Although electronic structure calculations at zero temperature can routinely handle systems with hundreds to even thousands of atoms, accounting for finite-temperature effects becomes crucial in many applications. Such calculations must therefore be set in a statistical ensemble, where the leading-order entropic contribution arises from nuclear motion due to the typically large gaps in electronic energies. In this context, because of the large mass of nuclei with respect to that of electrons, it is physically relevant to maintain electrons at the quantum level and to treat nuclei as classical particles. Nevertheless, the forces acting on the nuclei are computed from the solutions of the quantum electronic ground state, i.e. from the solutions of the Born Oppenheimer Hamiltonian. This forms the basis of ab-initio molecular dynamics Iftimie, Minary, and Tuckerman (2005); Silvestrelli and Parrinello (1999) which recovers the classical nuclear entropic contributions through either classical Monte Carlo algorithms or classical trajectories under the ergotic hypothesis.
Nevertheless, it is important to keep in mind that nuclear quantum effects play a significant role for light nuclei and can substantially alter the rates of certain chemical processes, such as proton transfer reactionsBała et al. (1996); Hanna and Kapral (2005); Kim and Hammes-Schiffer (2006); Costentin, Robert, and Savéant (2010); Hammes-Schiffer and Stuchebrukhov (2010). Moreover, such proton transfers can also be strongly influenced by the solvation effects, specially when polar solvents are usedKapral (2015). In this context, two fundamentally different yet interconnected nuclear quantum effects are at work: (i) the intrinsic delocalization of nuclear wave packets due to the quantum nature of nuclei, and (ii) the nonadiabatic couplings that arise between various adiabatic electronic states. Nuclear delocalization can be captured through explicit nuclear wave functions (see, for example, Ref. Beck et al., 2000 for a review of time-dependent mean-field approaches, and Ref. Hammes-Schiffer, 2021 for the nuclear-electron orbital approach) or via path-integral formalisms such as ring-polymer molecular dynamicsHabershon et al. (2013). Nonadiabatic couplings, on the other hand, are commonly addressed using surface hopping methods based on classical nuclear trajectoriesTully (1990); Hammes-Schiffer and Tully (1994); Ben-Nun and Martı’nez (2000); Li et al. (2005); Richardson and Thoss (2013); Ananth (2013), modified time-dependent mean-field techniquesWorth et al. (2008), semiclassical expansions of the density matrix and its Wigner representationWorth et al. (2008), or even hydrodynamic formulations of quantum mechanics à la Bohm-de BroglieBurghardt and Cederbaum (2001a, b); Burghardt and Bagchi (2006). In contrast, the present work adopts a fully classical treatment of the solvent molecules and focuses on clarifying the formally exact framework for a given QM/MM partitioning, as typically addressed in ab initio molecular dynamics.
Despite its considerable achievements, ab-initio molecular dynamics is computationally demanding due to the extensive number of QM calculations required to sample the exponentially large phase space. The convergence rate of the statistics can also vary significantly depending on the targeted properties and conditionsPan et al. (2013); Rozsa et al. (2018). To alleviate ab-initio molecular dynamics computational burden, QM/MMWarshel and Levitt (1976); Mulholland, Lyne, and Karplus (2000); Senn and Thiel (2009a); Groenhof (2013) methods were developed, where only a small, critical portion of the system is treated at the QM level, while the remainder is handled at the classical level (MM). The interaction between the classical particles are modeled using parametrized force fields and their dynamics obey Newton’s law of motion. The choice of the classical model to represent molecular entities together with the choice of the QM/MM interaction lacks of a unique definition and consists into a whole field of researchRiniker (2018); Harrison et al. (2018); Unke et al. (2021); Wu et al. (2023); Böselt, Thürlemann, and Riniker (2021); Ho et al. (2025) involving both physically-motivated and machine-learning oriented approaches. However, despite the significant gain in computational cost with respect to ab initio molecular dynamics, sampling the phase space of large systems remains computationally intensive even in QM/MM approaches.
Inspired by the extension of eDFT to QM systems at finite temperaturesMermin (1965), density-based formalisms have emerged for describing purely classical objects in statistical mechanics, particularly in the form of classical density functional theory (cDFT)Evans (1979). Originally rooted in liquid state theory, cDFT is essentially a classical counterpart to eDFT, reformulating the classical statistical mechanics problem of large ensembles of identical particles (atoms or molecules) in terms of one-body density functionals. cDFT, which is naturally formulated within the grand canonical ensemble, establishes the existence of a unique functional of the classical particle density that reaches its minimum at the equilibrium density. At this minimum, the functional is equal to the grand potential, which serves as the thermodynamic potential of the grand canonical ensemble. Therefore, most equilibrium properties can be obtained through functional minimization over the classical one-body density rather than through tedious sampling of configurations in an exponentially large phase space.
The efficiency of both eDFT and cDFT stems from a common principle: each method replaces the challenge of computing a high-dimensional object, such as the wavefunction in QM or the probability distribution in MM, with the optimization of a functional that depends only on a low-dimensional quantity, i.e. the one-body density of either quantum electrons or classical molecules. The mathematical framework allowing for such a simplification is the constrained-search principle introduced by Levy and LiebLevy (1979); Lieb (1983) which enables, under certain constraints of representativity, to give an explicit expression of the universal functionals of eDFT and cDFTDwandaru and Schmidt (2011), although not usable in practical calculations.
Its computational efficiency makes cDFT appealing to model solvent effect at the classical level while retaining a quantum description of the soluteTang et al. (2020); Petrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024). In this approach, the total Hamiltonian is partitioned into well defined QM and MM Hamiltonians. However, just as in molecular dynamics based QM/MM approaches, the coupling interaction between the QM and MM regions lacks of a unique practical definition. The coupling Hamiltonian is usually described as the sum of an electrostatic interaction combined with a non-bonded term which is often modeled using Lennard-Jones force field to account for repulsion and dispersion effectsSenn and Thiel (2009b); Giovannini, Lafiosca, and Cappelli (2017); Clemente, Capece, and Martí (2023). The Mixed cDFT-eDFT approach advantageously replaces the exhaustive phase-space sampling required in QM/MM molecular dynamics with a self-consistent optimization of both classical and quantum functionals. While the idea of merging cDFT and quantum methods is completely natural its theoretical justification is not immediately straightforward.
Petrosyan and co-workersPetrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007) introduced the joint density functional theory which consists in a DFT formulation of the QM/MM problem. The authors obtain a multi component electron/nuclei/solvent DFT formulation, which is justified by the use of a Levy-Lieb constrained search procedure allowing to express the universal functional as depending on the electron density and the nuclear density rather than the full density matrix describing the electrons and nuclei. However, to derive a QM/MM formulation, they treat the environment particles as classical by ’integrating out’ the electronic density associated with these nuclei, without explicitly indicating how this is done. Therefore, this step makes it difficult to pursue a rigorous derivation. Their working functional is instead defined as the sum of the Quantum Kohn-Sham functional describing the electrons of the solute, the classical functional describing the solvent and a coupling functional defined as the difference between the exact functional and the sum of the two former functionals. In other workJeanmairet, Levesque, and Borgis (2020); Tang et al. (2020), a similar form of the functional was invoked without further theoretical justification.
Thus, while QM/cDFT approaches offer potentially attractive computational scaling and appears as a quite natural outcome, their derivation from an exact theory remains somewhat ambiguous. This work addresses this gap by establishing a rigorous mathematical framework that provides a DFT-equivalent description of mixed QM/MM systems. Here, we focus on the canonical ensemble, deriving equations to compute the Helmholtz free energy using only the QM and MM one-body densities. Extending this formulation to the grand canonical ensemble for classical particles is conceptually straightforward but introduces additional complexity in the notations due to the varying number of particles. We therefore do not include it in the core of the paper, but rather in the supplementary materials. This derivation follows a similar route to its formally lighter canonical counterpart, which is reported in the present paper. The main difference between the two ensembles is the existence of a unique mapping between the equilibrium density and the external potential in the grand canonical ensemble. In contrast, in the canonical ensemble, the equilibrium density is associated to a family of external potentials, defined up to an additive constantGonzález et al. (1997); Lutsko (2022).
Our derivation begins with two key assumptions: (a) the entire system is partitioned into a QM region and a MM region; and (b) specific models are available for both the MM subsystem and the QM/MM coupling. As such, we do not address the QM/MM crossover or the accuracy of the MM and QM/MM models themselves. We assume these models are given, as is standard in practical QM/MM calculations.
Under these assumptions, the primary objective of this paper is to derive, in a straightforward manner, the one-body density functional theory (DFT) formulation for a QM/MM system. Therefore, The key steps of this derivation are as follows: i) We begin by defining the general Hamiltonian of the system, considering a broadly applicable QM/MM Hamiltonian (see Sec. II.1). ii) We then express the exact solution of the problem using full-system quantities (i.e., those describing all particles). This includes the Helmholtz free energy of the QM/MM system, along with the corresponding equilibrium density matrix and partition function (see Sec. II.2). iii) Next, we reformulate the problem as the minimization of a functional over generic full-system density matrices. In our case, this means expressing the Helmholtz free energy as a functional of QM/MM density matrices that describe all particles in the system (see Sec. II.3.1). iv) We then decompose this functional into intrinsic (arising from kinetic energy, inter-particle interactions and entropic contributions) and external (arising from external potentials) contributions (see Sec II.3.2). v) Under certain assumptions about the form of the external potentials, we show how the external part of the functional can be written as a linear functional of one-body densities. In our case, this results in a functional of both classical and quantum one-body densities (see Sec. III.1). vi) Leveraging the many-to-one correspondence between full-system and one-body quantities, we perform a Levy–Lieb constrained search to rewrite the intrinsic free energy functional as a universal functional of the one-body densities (see Sec. III.1). vii) We then establish a connection between electronic DFT (eDFT) and classical DFT (cDFT) by introducing a universal QM/MM correlation functional of the one-body QM and MM densities (see Sec. III.2). viii) Finally, we discuss this formulation in the context of solvation problems and recent developments in QM/cDFT coupling approaches (see Sec. III.3). The Appendix Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ contains a proof of the variational properties of the QM/MM functional introduced here, and the Supplementary Material provides a recall of the Von-Neuman variational principle in statistical QM for the canonical ensemble, an alternative derivation of the QM/MM equilibrium density previously established in the adiabatic basis by Nielsen et. al. in the context of QM/MM dynamicsNielsen, Kapral, and Ciccotti (2001), and an extension to the semi grand canonical ensemble of the present formalism. As this work is intended for both the QM and MM communities, which often differ in terminology and conceptual frameworks, we highlight certain elements throughout the text to aid accessibility. Some of these clarifications may appear self-evident to some readers; we hope, however, that they do not unduly burden the readability of the article.
II Quantum-classical system in the Canonical ensemble: equilibrium density and variational formulation
The objective of this section is to give a variational formulation for the Helmholtz free energy of a mixed QM/MM system based on a functional minimization over general densities involving $N_{{\text{qm}}}$ and $N_{{\text{mm}}}$ variables. This is essentially a QM/MM equivalent to the Gibbs variational principle in classical statistical physics or the Von-Neumann variational principle in quantum statistical physics (see the supplementary information for a short review). This variational principle, although hardly usable in practice due to the high-dimensionality of the objects it manipulates, is an important milestone to properly obtain the one-body formulation, which is the aim of this paper.
We define in Sec. II.1 the working Hamiltonian that describes the QM/MM system. Again, we do not discuss the validity of the partitioning or the quality of the coupling term, we assume that these preliminary considerations have already been addressed. The equilibrium density of a mixed QM/MM system is then given in Sec. II.2, and we briefly discuss the physical meaning of that result. We eventually show in Sec. II.3 that the Helmholtz free energy can be obtained through a variational principle over mixed QM/MM density matrices.
II.1 The Hamiltonian of the quantum-classical system
We consider a QM/MM system, where the partition and the choice of the model to describe the MM part and the coupling term have already been made
We adopt the following notation conventions: (i) lowercase letters denote purely quantum quantities (variables or operators), while uppercase letters denote classical or mixed quantum–classical quantities; (ii) bold symbols denote collective quantities (over all particles), while non-bold symbols denote single-particle or pairwise quantities. The only exception is the number of quantum particles, labeled $N_{{\text{qm}}}$. Atomic units are used throughout.
The total system is divided into a set of $N_{{\text{mm}}}$ identical classical particles, described by their classical positions ${\bf{Q}}=\{{{Q}}_{1},\ldots,{{Q}}_{N_{{\text{mm}}}}\}$ and linear momentum ${\bf P}=\{{{P}}_{1},\ldots,{{P}}_{N_{{\text{mm}}}}\}$ and a set of $N_{{\text{qm}}}$ identical quantum particles described by their position variables ${\bf{q}}=\{{{q}}_{1},\ldots,{{q}}_{N_{{\text{qm}}}}\}$.
The total Hamiltonian is written as the sum of the QM and MM Hamiltonians, and a QM/MM coupling term, namely
$\hat{\bf H}_{\text{tot}}({\bf{Q}},{\bf P},{\bf{q}})=\hat{\bf h}_{{\text{qm}}}({\bf{q}})+{\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}},{\bf{q}}).$ (1)
In the rest of the paper, explicit dependency on the QM and MM variables will be dropped for concision when unambiguous. In Eq.(1), $\hat{\bf h}_{{\text{qm}}}$ is the quantum Hamiltonian operator describing $N_{{\text{qm}}}$ quantum identical particles of mass $m$ (typically electrons). These particles interact with an external potential $v^{\text{ext}}_{{\text{qm}}}({{q}})$, which is local in the position representation, and with each other through a pairwise potential ${w}_{{\text{qm}}}({{q}}_{i},{{q}}_{j})$. The explicit form of $\hat{\bf h}_{{\text{qm}}}$ reads then
$\hat{\bf h}_{{\text{qm}}}=\hat{\bf t}+\hat{{\bf v}}^{\text{ext}}_{{\text{qm}}}+\hat{\bf w}_{{\text{qm}}},$ (2)
with
$\hat{\bf t}=\sum_{i=1}^{N_{{\text{qm}}}}\frac{\hat{p}_{i}^{2}}{2m},\qquad\hat{p}_{i}=-i\nabla_{{{q}}_{i}},$ (3)
$\hat{{\bf v}}^{\text{ext}}_{{\text{qm}}}=\sum_{i=1}^{N_{{\text{qm}}}}v^{\text{ext}}_{{\text{qm}}}({{q}}_{i}),\qquad\hat{\bf w}_{{\text{qm}}}=\sum_{j=1}^{N_{{\text{qm}}}}\sum_{i>j}{w}_{{\text{qm}}}({{q}}_{i},{{q}}_{j}).$ (4)
Then, ${\bf H}_{{\text{mm}}}$ is the classical model Hamiltonian function representing $N_{{\text{mm}}}$ identical particles of mass $M$, in interaction with a local external potential $V^{\text{ext}}_{{\text{mm}}}({{Q}})$ and interacting among themselves through a local pairwise potential ${W}_{{\text{mm}}}({{Q}}_{i},{{Q}}_{j})$. These are typically effective particles, such as atoms or rigid molecules, and they interact with each other through parametrized force fields. The explicit form of the classical Hamiltonian ${\bf H}_{{\text{mm}}}$ reads then
${\bf H}_{{\text{mm}}}={\bf T}+{\bf V}^{\text{ext}}_{{\text{mm}}}+{\bf W}_{{\text{mm}}},$ (5)
with
${\bf T}=\sum_{i=1}^{N_{{\text{mm}}}}\frac{{{P}}_{i}^{2}}{2M},\qquad{\bf V}^{\text{ext}}_{{\text{mm}}}=\sum_{i=1}^{N_{{\text{mm}}}}V^{\text{ext}}_{{\text{mm}}}({{Q}}_{i}),$ (6)
${\bf W}_{{\text{mm}}}=\sum_{j=1}^{N_{{\text{mm}}}}\sum_{i>j}{W}_{{\text{mm}}}({{Q}}_{i},{{Q}}_{j}).$ (7)
Regarding the coupling term ${W}_{{\text{qm}}}^{{\text{mm}}}$, we simply assume here that it is an additive pairwise potential, i.e.
$\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}=\sum_{i=1}^{N_{{\text{qm}}}}\sum_{j=1}^{N_{{\text{mm}}}}{W}_{{\text{qm}}}^{{\text{mm}}}({{q}}_{i},{{Q}}_{j})$ (8)
The function ${W}_{{\text{qm}}}^{{\text{mm}}}({q},{Q})$ represents the effective interaction between an effective particle (such as a nuclei, an atom or a molecule) at position Q{Q} and a fundamental particle (such as an electron) at position q{q}.
II.2 Mixed quantum-classical equilibrium density, partition function and Helmholtz free energy
We start by defining the exact equilibrium density matrix (EDM) of the QM/MM system in the canonical ensemble. This quantity is the analogue of the equilibrium density matrix of a purely quantum system or, equivalently, of the equilibrium probability density of a purely classical (MM) system.
With the Hamiltonian of Eq. (1), the EDM takes the following form
$\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})$ $=\frac{{\langle{\bf{q}}|}e^{-\beta\hat{\bf H}_{\text{tot}}({\bf{Q}},{\bf P})}{|{\bf{q}}^{\prime}\rangle}}{Z^{\text{tot}}}$ (9)
$=\frac{{\langle{\bf{q}}|}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}{|{\bf{q}}^{\prime}\rangle}e^{-\beta{\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})}}{Z^{\text{tot}}}.$
In Eq. (9) $\beta=T^{-1}$ in atomic units, and $|{\bf{q}}\rangle$} and $|{\bf{q}}^{\prime}\rangle$} are position eigenstates for the $N_{{\text{qm}}}$ quantum variables. The dependency of $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}$ on the classical variables, $\bf{Q}$}, has been made explicit. Because Eq. (9) involves the exponential of a quantum operator, the QM/MM EDM is itself a quantum operator, namely a quantum density matrix (i.e. positive definite and of trace unity). However, it is parametrized by the classical variables $({\bf{Q}},{\bf P})$ and therefore the QM system depends on the configuration of the classical particles. To understand the physical meaning of the EDM, it is instructive to analyse its diagonal elements, i.e. when ${\bf{q}}={\bf{q}}^{\prime}$. The quantity $\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}})$ gives the probability of finding the quantum particles at positions ${\bf{q}}=\{{{q}}_{1},\ldots,{{q}}_{N_{{\text{qm}}}}\}$ while the classical particles have positions ${\bf{Q}}=\{{{Q}}_{1},\ldots,{{Q}}_{N_{{\text{mm}}}}\}$ and momenta ${\bf P}=\{{{P}}_{1},\ldots,{{P}}_{N_{{\text{mm}}}}\}$. Off-diagonal elements encode purely quantum effects, such as nonlocality, and are necessary to compute the action of quantum operators—for example, the kinetic energy operator, which involves a Laplacian in the quantum subsystem.
A pictorial illustration of the physical meaning of the equilibrium density of Eq. (9) is provided by the Born-Oppenheimer description of a molecule, i.e. QM electrons and classical nuclei. In that case, $\hat{\bf h}_{{\text{qm}}}$ is composed of the kinetic operator of the electrons and their mutual coulomb interaction, with no external potential for either the QM or MM parts. The operator $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$ represents the nuclei-electron attraction, while ${\bf W}_{{\text{mm}}}({\bf{Q}})$ represents the Coulomb repulsion between the nuclei. Thus, $\hat{\bf H}_{\text{tot}}({\bf{Q}},{\bf P})$ entering in the exponential of Eq. (9) corresponds the Born-Oppenheimer Hamiltonian.
From a more formal point of view, it is worth noticing that Eq. (9) corresponds to the real-space representation of the EDM, originally obtained by Nielsen et. al.Nielsen, Kapral, and Ciccotti (2001) in the adiabatic basis, i.e. in the basis of eigenvectors of $\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$. In this basis, the density matrix is diagonal , as found by Nielsen et. al.Nielsen, Kapral, and Ciccotti (2001). In the case of quantum electrons and classical nuclei, the adiabatic states are the eigenstates of the Born-Oppenheimer Hamiltonian. Therefore, Eq. (9) ignores the non-adiabatic couplings between these states. In the context of ab initio MD, this implies that $\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})$ can be obtained from classical nuclear trajectories, with forces averaged over all adiabatic electronic states weighted by the corresponding Boltzmann factors. Consequently, it entirely ignores surface hopping between different adiabatic states, which is why $\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})$ is referred to as the adiabatic QM/MM density matrix. There is a flourishing literature addressing methods going beyond this adiabatic approximation, specially in the context of quantum dynamics or statistical mechanics for QM/MM systems (see for instance Ref. Kapral, 2015 for a review), but treatment of non adiabatic couplings is beyond the scope of this paper.
[…]
In Eq. (9), the normalization corresponds to the partition function of the QM/MM system which is expressed as
$Z^{\text{tot}}=\frac{1}{N_{{\text{mm}}}!}$ $\int\text{d}{\bf{Q}}\text{d}{\bf P}\,e^{-\beta{\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})}$ (10)
$\int\!\text{d}{\bf{q}}{\langle{\bf{q}}|}{e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}}{|{\bf{q}}\rangle}.$
The corresponding Helmholtz free energy is therefore given by
$F_{0}^{\text{tot}}=-T\log(Z^{\text{tot}}).$ (11)
Any equilibrium property $\langle O\rangle_{\text{eq}}$ can be obtained from derivatives of $Z^{\text{tot}}$. Alternatively, $\langle O\rangle_{\text{eq}}$ can be computed as the trace of the product of $\mathcal{M}$ and the operator $\hat{O}$, which may depend on both quantum and classical variables
$\langle O\rangle_{\text{eq}}$ $=\int\text{d}{\bf{Q}}\text{d}{\bf P}\,\int\!\text{d}{\bf{q}}\text{d}{\bf{q}}^{\prime}\hat{O}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}}^{\prime},{\bf{q}})$ (12)
$\equiv\text{Tr}\{\hat{O}\hat{\mathcal{M}}\}$.
Although the expression of the equilibrium density in Eq. (9) is relatively simple, it benefits from further comments. Firstly, as usual in classical statistical mechanics, the contribution to the Helmholtz free energy coming from the purely classical kinetic term in ${\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})$ can be computed analytically. Therefore, it is convenient to separate the equilibrium density according to its dependence on $\bf{Q}$} and $\bf P$}
$\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})=\frac{e^{-\beta{\bf T}}}{z_{{\bf P}}}\mathcal{P}_{\text{eq}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime}),$ (13)
where $\bf T$} is the purely classical kinetic energy defined in Eq. (6), and
$\mathcal{P}_{\text{eq}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})=\frac{{\langle{\bf{q}}|}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}{|{\bf{q}}^{\prime}\rangle}e^{-\beta{{\bf{V}}}_{\text{mm}}({\bf{Q}})}}{Z},$ (14)
${{\bf{V}}}_{\text{mm}}$ $={\bf V}^{\text{ext}}_{{\text{mm}}}+{\bf W}_{{\text{mm}}},$ (15)
$z_{{\bf P}}$ $=\frac{1}{N_{{\text{mm}}}!}\int\text{d}{\bf P}e^{-\beta{\bf T}}=\frac{1}{N_{{\text{mm}}}!}\bigg(\frac{2\pi M}{\beta}\bigg)^{\frac{3}{2}N_{{\text{mm}}}},$ (16)
$Z$ $=\int\text{d}{\bf{Q}}\int\text{d}{\bf{q}}{\langle{\bf{q}}|}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}{|{\bf{q}}\rangle}e^{-\beta{{\bf{V}}}_{\text{mm}}({\bf{Q}})},$ (17)
$Z^{\text{tot}}$ $=z_{{\bf P}}Z,$ (18)
such that the free energy can be written as
$F_{0}^{\text{tot}}$ $=F_{0}+F_{{\bf P}},$ (19)
where
$F_{{\bf P}}=-T\log(z_{{\bf P}}),$ (20)
and
$F_{0}=-T\log(Z).$ (21)
$F_{0}$ represents the contribution to the Helmholtz free energy of the QM/MM system arising from particle interactions. Computing $F_{0}$ from Eq. (21) is infeasible for realistic systems, as it requires evaluating Eq. (17). The aim of this paper is to develop an alternative, more practical, route to $F_{0}$.
The EDM $\mathcal{P}_{\text{eq}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})$ of Eq.(14) can be decomposed into the product of a purely quantum density matrix ${\boldsymbol{\rho}}_{\text{eq}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})$ (more precisely its matrix element evaluated in real space) and a purely classical function $\boldsymbol{p}_{\text{eq}}({\bf{Q}})$, i.e.
$\mathcal{P}_{\text{eq}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})={\boldsymbol{\rho}}_{\text{eq}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})\boldsymbol{p}_{\text{eq}}({\bf{Q}}),$ (22)
with
${\boldsymbol{\rho}}_{\text{eq}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})$ $={\langle{\bf{q}}|}\hat{{\boldsymbol{\rho}}\,}_{\text{eq}}({\bf{Q}}){|{\bf{q}}^{\prime}\rangle},$ (23)
$\hat{{\boldsymbol{\rho}}\,}_{\text{eq}}({\bf{Q}})=\frac{1}{Z^{{\text{qm}}}({\bf{Q}})}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))},$ (24)
$Z^{{\text{qm}}}({\bf{Q}})$ $=\int\text{d}{\bf{q}}{\langle{\bf{q}}|}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}{|{\bf{q}}\rangle},$ (25)
$\boldsymbol{p}_{\text{eq}}({\bf{Q}})=\frac{Z^{{\text{qm}}}({\bf{Q}})}{Z}e^{-\beta{{\bf{V}}}_{\text{mm}}({\bf{Q}})}.$ (26)
$Z$ and $Z^{{\text{qm}}}({\bf{Q}})$ are related through the following equation
$Z$ $=\int\text{d}{\bf{Q}}Z^{{\text{qm}}}({\bf{Q}})e^{-\beta{{\bf{V}}}_{\text{mm}}({\bf{Q}})}.$ (27)
It is noteworthy that, for each value of $\bf{Q}$}, $\hat{{\boldsymbol{\rho}}\,}_{\text{eq}}({\bf{Q}})$ defined in Eq. (23) and (24) is a density matrix, i.e. positive definite and of unity norm. Similarly, $\boldsymbol{p}_{\text{eq}}$ of Eq. (26) is a probability distribution. Therefore, $\hat{{\boldsymbol{\rho}}\,}_{\text{eq}}$ and $\boldsymbol{p}_{\text{eq}}$ are special cases of more general objects: density matrices representing $N_{{\text{qm}}}$ quantum particles and probability distribution functions representing $N_{{\text{mm}}}$ classical particles, respectively. This will allow to formulate a variational formulation of the Helmholtz free energy for a QM/MM system in Sec. II.3.
Last but not least, the decomposition of the equilibrium density of Eq. (22) into a product of a purely quantum equilibrium density and a purely classical equilibrium density might suggest that the quantum and classical subsystems are uncorrelated, even though they interact. Nevertheless, the quantum part of the equilibrium density, $\hat{{\boldsymbol{\rho}}\,}_{\text{eq}}$, depends on the coordinates of the classical systems through the QM/MM interaction potential $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$. The classical part, $\boldsymbol{p}_{\text{eq}}$, is in turn affected by the quantum degrees of freedom through its normalization, $Z^{{\text{qm}}}({\bf{Q}})$. To see this more explicitly, one can notice that the form of Eq. (26) can be rewritten as
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{p}_{\text{eq}}({\bf{Q}})=\frac{1}{Z}e^{-\beta\big({{\bf{V}}}_{\text{mm}}({\bf{Q}})+U_{\text{eff}}({\bf{Q}})\big)},}$ (28)
where now the classical particles experience an additional the effective potential $U_{\text{eff}}({\bf{Q}})$ due to the presence of the QM system, which is defined as
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}U_{\text{eff}}({\bf{Q}})=-\frac{1}{\beta}\log\big(Z^{{\text{qm}}}({\bf{Q}})\big).$} (29)
Therefore, the two systems are correlated.
II.3 Variational formulation of the Helmholtz free energy for a QM/MM system
Both classical and quantum DFT can be formulated through the Levy-Lieb constrained-search formalism. The first step is to express the quantum ground state energy and the classical Helmholtz free energy as a general functional minimization problem, defined over the many-body functions, namely the electronic density matrix and the probability distribution, respectively.
Since the goal of this work is to obtain a DFT formulation of the QM/MM problem, the first step is to reformulate the Helmholtz free energy of Eq. (21) as a functional minimization problem over QM/MM density matrices whose form generalize the EDM of Eq. (22). This is the aim of this section. To the best of our knowledge, this was never formalized before.
II.3.1 Variational formulation
As we know that the equilibrium density for a mixed QM/MM system is expressed as a special case of product of a quantum density matrix by a classical probability distribution, we can now reformulate the Helmholtz free energy $F_{0}$ as a variational problem over QM and MM objects, such that the equilibrium density of Eq. (22) is the minimizer of that functional and that it yields the correct Helmholtz free energy of Eq. (21).
Thus, we would like to write the Helmholtz free energy as the minimum of a functional defined on a couple of quantum and classical objects as follows
$F_{0}=\min_{\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}}F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}],$ (30)
where the operator $\hat{{\boldsymbol{\rho}}\,}$ is a quantum density matrix representing $N_{{\text{qm}}}$ particles but also parametrically depending on all classical variables ${\bf{Q}}=\{{{Q}}_{1},\ldots,{{Q}}_{N_{{\text{mm}}}}\}$, i.e.
$\hat{{\boldsymbol{\rho}}\,}\equiv\hat{{\boldsymbol{\rho}}\,}({\bf{Q}}),$ (31)
and $\boldsymbol{p}$ is a classical probability density representing $N_{{\text{mm}}}$ classical particles and not depending on any quantum variable. Therefore, $\hat{{\boldsymbol{\rho}}\,}$ and $\boldsymbol{p}$ must be normalized to unity in order to be acceptable density matrices and probability measures, respectively, i.e
$\int\text{d}{\bf{q}}\,\,{\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})=1\quad\forall\,\,{\bf{Q}},$ $\int\text{d}{\bf{Q}}\,\,\boldsymbol{p}({\bf{Q}})=1,$ (32)
where ${\boldsymbol{\rho}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})$ is the real-space representation of the operator $\hat{{\boldsymbol{\rho}}\,}$, i.e.
${\boldsymbol{\rho}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})={\langle{\bf{q}}|}\hat{{\boldsymbol{\rho}}\,}({\bf{Q}}){|{\bf{q}}^{\prime}\rangle},$ (33)
such that that the product of the two functions is normalized to unity
$\int\text{d}{\bf{Q}}\int\text{d}{\bf{q}}\,\,{\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})\boldsymbol{p}({\bf{Q}})=1.$ (34)
Therefore, the equilibrium density of the mixed QM/MM system described in Eq. (22) lies within the functional space defined by all the couple of functions and density matrices satisfying Eqs. (32)-(34). As shown in the appendix (see section Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$), if we define the Helmholtz functional as follows
$F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]=\mathcal{U}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]-T\mathcal{S}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}],$ (35)
where $\mathcal{U}$ is an energy functional excluding the classical kinetic term (which has already been accounted for in $F_{{\bf P}}$)
$\mathcal{U}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]=\text{Tr}\big\{\hat{{\boldsymbol{\rho}}\,}\boldsymbol{p}(\hat{\bf h}_{{\text{qm}}}+\hat{{\bf v}}^{\text{ext}}_{{\text{qm}}}+{\bf W}_{{\text{mm}}}+{\bf V}^{\text{ext}}_{{\text{mm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}})\big\},$ (36)
and $\mathcal{S}$ is the dimensionless entropy of the total system
$\mathcal{S}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]=-\text{Tr}\big\{\hat{{\boldsymbol{\rho}}\,}\boldsymbol{p}\log\big(\hat{{\boldsymbol{\rho}}\,}\boldsymbol{p}\big)\big\},$ (37)
then the only minimizer of the functional of Eq. (35) is $(\hat{{\boldsymbol{\rho}}\,}_{\text{eq}},\boldsymbol{p}_{\text{eq}})$ of Eq. (22), i.e
$-T\log(Z)\leq F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]\quad\forall\,\,(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}),$ (38)
and
$F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]=F_{0}=-T\log(Z)\Leftrightarrow(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})=(\hat{{\boldsymbol{\rho}}\,}_{\text{eq}},\boldsymbol{p}_{\text{eq}}).$ (39)
This variational formulation, is the analogue of the Gibbs or Von Neumann variational principles, but applied to a QM/MM system. As with any variational formulation, it naturally enables the construction of approximations by imposing constraints on the variables of minimization, the archetypal example being the mean-field approach, which is briefly discussed in Sec. III.3.
II.3.2 Rewriting of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$
Because the function $\boldsymbol{p}$ does not explicitly depend on the quantum variables and because we chose the normalization conditions of Eq. (32), we can naturally decompose the entropy as
$\mathcal{S}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ $=S_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]+S_{{\text{mm}}}[\boldsymbol{p}],$ (40)
where $S_{{\text{mm}}}[\boldsymbol{p}]$ is the purely classical dimension-less entropy
$S_{{\text{mm}}}[\boldsymbol{p}]$ $=-\int\text{d}{\bf{Q}}\,\boldsymbol{p}({\bf{Q}})\log(\boldsymbol{p}({\bf{Q}})),$ (41)
and $S_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ is the quantum-classical dimension-less entropy
[…]
$=-\int\text{d}{\bf{Q}}\,\boldsymbol{p}({\bf{Q}})\int\text{d}{\bf{q}}\,{\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})\log({\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})),$
which depends explicitly on both $\hat{{\boldsymbol{\rho}}\,}$ and $\boldsymbol{p}$.
We can similarly decompose $\mathcal{U}$ as follows
$\mathcal{U}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ $=E_{{\text{mm}}}[\boldsymbol{p}]+E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ (43)
$+E_{\text{qm}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]+E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]+E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}],$
where $E_{\text{mm}}[\boldsymbol{p}]$ is the mutual interaction of the classical part
$E_{\text{mm}}[\boldsymbol{p}]$ $=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}}){\bf W}_{{\text{mm}}}({\bf{Q}}),$ (44)
$E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ is the functional accounting for the interaction of the classical particles with the external potential
$E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ $=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}}){\bf V}^{\text{ext}}_{{\text{mm}}}({\bf{Q}}),$ (45)
$E_{\text{qm}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ is the kinetic and mutual interaction of the quantum part
$E_{\text{qm}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ $=\text{Tr}\{(\hat{\bf t}+\hat{\bf w}_{{\text{qm}}})\boldsymbol{p}\hat{{\boldsymbol{\rho}}\,}\}$ (46)
$=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}})$
$\quad\int\text{d}{\bf{q}}\text{d}{\bf{q}}^{\prime}(\bf{t}({\bf{q}},{\bf{q}}^{\prime})+{w}_{{\text{qm}}}({\bf{q}},{\bf{q}}^{\prime})){\boldsymbol{\rho}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}}),$
$E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ is the functional accounting for the interaction of the quantum particles with the external potential
$E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ $=\text{Tr}\{\hat{{\bf v}}^{\text{ext}}_{{\text{qm}}}\boldsymbol{p}\hat{{\boldsymbol{\rho}}\,}\}$ (47)
$=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}})\int\text{d}{\bf{q}}{\bf v}^{\text{ext}}_{{\text{qm}}}({\bf{q}},{\bf{q}}){\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}}),$
and eventually $E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ is the quantum-classical interaction,
$E_{{\text{qm}}}^{{\text{mm}}}$ $[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]=\text{Tr}\{\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}\boldsymbol{p}\hat{{\boldsymbol{\rho}}\,}\}$ (48)
$=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}})\int\text{d}{\bf{q}}\int\text{d}{\bf{q}}^{\prime}{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime}){\boldsymbol{\rho}}({\bf{q}}^{\prime},{\bf{q}},{\bf{Q}}).$
It is worth noting that the form of Eq. (48) is quite general, since ${\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})$ is a priori non local. This form of interaction includes, for example, Lennard-Jones potentials, frozen density Coulomb interaction or density overlap models for Pauli repulsionWheatley and Price (1990).
Eventually, once all terms are inserted in Eq. (30), the Helmholtz free energy $F_{0}$ can be obtained as
$F_{0}=\min_{\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}}\big\{$ $E_{\text{qm}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]+E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]+E_{{\text{mm}}}[\boldsymbol{p}]+E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ (49)
$+E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]-T\big(S_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]+S_{{\text{mm}}}[\boldsymbol{p}]\big)\big\}.$
Eq. (49) therefore provides a variational formulation of $F_{0}$ for a mixed QM/MM system, expressed as a minimization over the couple $(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})$. One can see from Eqs. (41), (44) and (45) that the functionals corresponding to the MM part of the energy and entropy (i.e. $E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$, $E_{{\text{mm}}}[\boldsymbol{p}]$ and $S_{{\text{mm}}}[\boldsymbol{p}]$) are, per se, completely independent of the quantum part of the system. They only depend on the classical probability $\boldsymbol{p}$. Nevertheless, the quantum and classical parts remain coupled in the minimization of Eq. (49) through the coupling functional $E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$. In contrast, the QM parts energy and entropy functionals explicitly depend on the classical MM probability distribution $\boldsymbol{p}$.
Although Eq. (49) provides a variational formulation of the Helmholtz free energy for a QM/MM system, the minimization variables $(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})$ remain very high-dimensional because they explicitly describe every individual particle in the system. The objective of the following section is to reformulate the variational principle of Eq. (49) using functionals depending on much lower-dimensional objects, i.e. the one-body QM and MM densities.
III One-body density formulation through Levy-Lieb constrained search formulation
Starting from the variational formulation of the Helmholtz free energy for the QM/MM system in terms of the $N_{{\text{qm}}}+N_{{\text{mm}}}$ quantities in Eq. (49), we want to obtain a one-body DFT formulation. As in classical and quantum DFT, the key idea is to split the total energy into an external part and an intrinsic part. More precisely, because of the one-body nature of both classical and quantum external potentials $V^{\text{ext}}_{{\text{mm}}}({Q})$ and $v^{\text{ext}}_{{\text{qm}}}({q})$, and the permutational symmetry of both $\boldsymbol{p}$ and $\hat{{\boldsymbol{\rho}}\,}$, the corresponding energy terms $E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ and $E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ in Eqs. (45) and (47) can be expressed in terms of the one-body densities. Then, by applying the Levy-Lieb constrained search, we can also reformulate the intrinsic part of the energy as a functional of the one-body quantum and classical densities. Therefore, the variational problem of Eq. (49) is directly expressed as a functional of the one-body QM and MM densities rather than in terms of full $N_{{\text{mm}}}+N_{{\text{qm}}}$ variables quantities.
III.1 Classical and quantum one-body density and the Levy-Lieb constrained search
The classical one-body density is naturally defined as the trace over all quantum variables and all classical variables but one of the product $\hat{{\boldsymbol{\rho}}\,}({\bf{q}},{\bf{Q}})\boldsymbol{p}({\bf{Q}})$, which, thanks to the normalization of Eq. (32), is simply
[…]
Similarly we can define the QM one-body density $\rho({q})$ as
[…]
Thanks to the permutational symmetry of both $\hat{{\boldsymbol{\rho}}\,}({\bf{Q}})$ and $\boldsymbol{p}({\bf{Q}})$ with respect to the exchange of either QM or MM coordinates, the two external functionals can be rewritten as functionals of the QM and MM one-body densities, respectively
$E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ $=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}}){\bf V}^{\text{ext}}_{{\text{mm}}}({\bf{Q}})$ (52)
$=\int\text{d}{Q}n({Q})V^{\text{ext}}_{{\text{mm}}}({Q})$
$\equiv(V^{\text{ext}}_{{\text{mm}}}|n),$
$E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ $=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}})\int\text{d}{\bf{q}}{\bf v}^{\text{ext}}_{{\text{qm}}}({\bf{q}},{\bf{q}}){\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})$ (53)
$=\int\text{d}{q}\rho({q})v^{\text{ext}}_{{\text{qm}}}({q})$
$\equiv(v^{\text{ext}}_{{\text{qm}}}|\rho).$
Since the contributions arising from the external potentials can be directly expressed in terms of the one-body densities $(\rho,n)$, we now introduce a universal functional of $(\rho,n)$ which contains all the information required to compute the free energy. This functional is built using the Levy-Lieb constrained search formalism, as in eDFT. First, we introduce the intrinsic Helmholtz free energy functional of the $N$-body densities
$f[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ $=E_{{\text{mm}}}[\boldsymbol{p}]+E_{\text{qm}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]+E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]-T\mathcal{S}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}].$ (54)
such that we can then rewrite the Helmholtz free energy as
$F_{0}=\min_{\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}}\big(E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]+E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]+f[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]\big).$ (55)
Then, the many-to-one correspondence between $(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})$ and $(\rho,n)$ allows us to introduce the Levy-Lieb constrained search as follows
$\min_{\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}}F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]=\min_{(\rho,n)}\,\,\min_{(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})\rightarrow(\rho,n)}F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}],$ (56)
where the notation $(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})\rightarrow(\rho,n)$ means that the search is performed over the set of couple of $N_{{\text{qm}}}+N_{{\text{mm}}}$-body densities $(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})$ that yield the 1-body densities $(\rho,n)$. Therefore, by remembering that the external parts depends only on the one-body densities (see Eqs. (52) and (53)), one can rewrite Eq. (56) as
$\min_{\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}}F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]=$ $\min_{(\rho,n)}\big[(V^{\text{ext}}_{{\text{mm}}}|n)+(v^{\text{ext}}_{{\text{qm}}}|\rho)$ (57)
$+$ $\min_{(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})\rightarrow(\rho,n)}f[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]\big].$
We have just introduced the universal Levy-Lieb functional of the one-body classical and quantum density,
$\mathcal{F}[\rho,n]=\min_{(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})\rightarrow(\rho,n)}f[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}],$ (58)
such that one can rewrite the Gibbs free energy as
$F_{0}=\min_{(\rho,n)}\big((V^{\text{ext}}_{{\text{mm}}}|n)+(v^{\text{ext}}_{{\text{qm}}}|\rho)+\mathcal{F}[\rho,n]\big).$ (59)
This provides a constructive proof of the existence of a variational principle for the Helmholtz Free energy, $F_{0}$, as a functional of one-body densities.
III.2 Explicit coupling in terms of eDFT and cDFT
Although Eq. (59) expresses the Helmholtz free energy only in terms of the one-body densities $(\rho,n)$, the functional itself remains unknown in practice. Thus, the explicit connection with well known universal quantum and classical density functionals is still unclear. This paragraph focuses on rewriting of $\mathcal{F}[\rho,n]$ so that the functionals of MerminMermin (1965) and EvansEvans (1979) explicitly appear, along with a new QM/MM correlation functional.
We begin by recalling the universal purely quantum one-body density functionals of Mermin
[…]
where $\hat{{\boldsymbol{\rho}}\,}$ is a quantum $N$-body density matrix and $(\hat{{\boldsymbol{\rho}}\,}\rightarrow\rho)$ means that the minimization is performed over $N$-body quantum density matrices $\hat{{\boldsymbol{\rho}}\,}$ leading to a given quantum one-body density $\rho$. Similarly, we can define the universal classical one-body functional of Evans
$\mathcal{F}_{\text{mm}}[n]=\min_{\boldsymbol{p}\rightarrow n}\text{Tr}\{\boldsymbol{p}\big({\bf W}_{{\text{mm}}}+T\log(\boldsymbol{p})\big)\},$ (61)
where $\boldsymbol{p}$ is a classical $N$-body probability distribution and $(\boldsymbol{p}\rightarrow n)$ means that the minimization is performed over $N$-body classical probability distribution $\boldsymbol{p}$ leading to a given classical one-body density $n$. Historically, both Mermin and Evans introduced their universal functionals in the Grand Canonical ensemble, so Eqs. (60) and (61) are actually the Canonical counter part of the original functionals. Also, in their respective original papersMermin (1965); Evans (1979), neither Mermin or Evans explicitly used the constrained search framework which was formalized later by Levy and LiebLevy (1979); Lieb (1983).
We can then rewrite $\mathcal{F}[\rho,n]$ as follows
$\mathcal{F}[\rho,n]=$ $\mathcal{F}_{\text{qm}}[\rho]+\mathcal{F}_{\text{mm}}[n]+\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]+\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n],$ (62)
where $\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ is the mean-field QM/MM interaction
$\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ $=\int\text{d}{q}\int\text{d}{Q}{W}_{{\text{qm}}}^{{\text{mm}}}({Q},{q})n({Q})\rho({q}),$ (63)
which is the QM/MM analogue of the Hartree-term in eDFT, and where $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ is the QM/MM correlation functional, i.e, the complementary functional
$\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]=\mathcal{F}[\rho,n]-\big(\mathcal{F}_{\text{qm}}[\rho]+\mathcal{F}_{\text{mm}}[n]+\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]\big).$ (64)
The functional $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ in Eq. (64) is there to reproduce all correlation terms arising because of the QM/MM interaction. The latter contains therefore some energetic and entropic terms.
We can then rewrite explicitly $F_{0}$ as
$F_{0}=\min_{(\rho,n)}\big\{$ $(v^{\text{ext}}_{{\text{qm}}}|\rho)+\mathcal{F}_{\text{qm}}[\rho]+\mathcal{F}_{\text{mm}}[n]+(V^{\text{ext}}_{{\text{mm}}}|n)$ (65)
$+\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]+\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]\big\},$
where we can explicitly recognize the purely quantum and purely classical free-energy functionals, together with the QM/MM coupling expressed as the sum of the mean-field and correlation terms. If $\delta\mathcal{F}_{{\text{qm}}}$ is ignored, we recover the mean-field coupling between the eDFT and cDFT functionals that is generally used in the literaturePetrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007); Tang et al. (2020); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024). We highlight here that, although the outcome of the present derivation is rather natural as it consists of a multi component DFT formalism, the explicit definition of $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ in Eq. (64) can help in future developments of approximations for $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$. For instance, the variational formulation in terms of full-body densities allows the use the adiabatic connection to reformulate the problem, and therefore potentially to import several successful strategies used in eDFT to approximate correlation functionals.
III.3 Applications to solvation problems
The development presented here were motivated by the necessity of a well grounded theoretical framework for a DFT description of a quantum solute solvated in a classical solvent. This section discusses the application of Eq. (65) to this particular problem.
As common in electronic DFT calculations with an implicit solvent description, the nuclei of the molecular solute are considered clamped. The Hamiltonian of the quantum system of Eq. (2) is therefore the Born-Oppenheimer Hamiltonian.
Regarding now the classical Hamiltonian, it strongly depends on the nature of the considered solvent. We simply assume here that the external part of this potential, ${\bf V}^{\text{ext}}_{{\text{mm}}}$, represents the interaction between the solvent and the classical nuclei of the solute, and that it is local and expressed as a sum of one-body terms. Therefore, ${\bf V}^{\text{ext}}_{{\text{mm}}}$ naturally contains a Coulombic term but additional terms can be incorporated, in order, for instance, to model repulsion and dispersion interactions. There exists multiple types of classical potential modeling dispersion and repulsion, the most popular one being the Lennard-Jones potential.
Central to the description of the QM/MM interaction, $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}$, accounts for the electron-solvent interaction. This is a part of the total Hamiltonian that is both unknown and whose quality is critical since it essentially accounts for the information that has been lost when replacing quantum particles by classical particles. This interaction is often modeled using a Coulomb potential. Since only the external potentials are required to be local to derive Eq. (65), more sophisticated models incorporating non-local quantum operators can also be incorporated in $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}$. This is a route that should be explored to improve the quality of the QM/MM model.
We now assume that, in the absence of the solvent, the typical electronic gap is large compared to the thermal agitation $T$, allowing us to neglect the contribution of electronic entropy in $\mathcal{F}_{\text{qm}}[\rho]$ (see Eq. (60)). This is equivalent to considering only the ground state of the electronic quantum system. To illustrate the validity of this approximation for systems at room temperature (T=300KT=300\,K), we note that the Boltzmann weight ratio between the ground state and an excited state with a gap of 0.25 eV ($\approx 0.01\,\text{Ha}$) is approximately $10^{-5}$. For a gap of 1.63 eV ($\approx 0.06\,\text{Ha}$), which marks the onset of the visible spectrum, this ratio drops dramatically to $10^{-28}$. Since most chemical systems near their equilibrium geometry exhibit electronic gaps of several eV, this approximation is well-suited for typical room-temperature conditions. As a consequence, we can take the limit $T\rightarrow 0$ in the definition of the Mermin universal functional (see Eq. (60)) to recover the usual universal Levy-Lieb electronic density functional, namely
[…]
Regarding solvent effects, we focus on cases where the solvent-solute interaction is not too strong, meaning there is no covalent bonding between the solute and solvent species. Under these conditions, the previous T=0KT=0~K assumption for $\mathcal{F}_{\text{qm}}[\rho]$ remains reasonable, as the electronic structure of the solute is not significantly altered by the presence of the solvent.
Previous attempts to apply mixed electronic-classical density functional theory to QM/MM systems were carried on in a semi-Grand canonical ensemble, where the number of classical particles is allowed to fluctuate. This choice was motivated by the availability of classical functionals, which are naturally formulated in the grand canonical ensemble. In the semi-Grand canonical ensemble, the state function is the grand potential, $\Omega_{0}$, for which the following variational principle, analogous to Eq. (65), stands
$\Omega_{0}=$ $\min_{(\rho,n)}\big\{\mathcal{F}_{\text{qm}}[\rho]+\Omega_{\text{mm}}[n^{\text{GC}}]+\delta\Omega_{{\text{qm}}}^{\text{mm}}[\rho,n^{\text{GC}}]$ (67)
$+\varepsilon_{{\text{qm}}}^{\text{mm}}[\rho,n^{\text{GC}}]+\left(v^{\text{ext}}_{{\text{qm}}}\mid\rho\right)+\left(V^{\text{ext}}_{{\text{mm}}}\mid n^{\text{GC}}\right)\big\}.$
In Eq. (67), $n^{\text{GC}}$ is the one-body density of classical particles, $\Omega_{\text{mm}}[n^{\text{GC}}]$ is the grand canonical universal functional a la Evans, $V^{\text{ext}}_{{\text{mm}}}$ is the external classical potential, $(v_{\text{ne}}|\rho)+\mathcal{F}_{\text{qm}}[\rho]$) is the usual quantum Born-Oppenheimer electronic functional and $\delta\Omega_{{\text{qm}}}^{\text{mm}}[\rho,n^{\text{GC}}]$ is the unknown QM/MM correlation term. Eq. (67) can be obtained following the same route that was used to derive Eq. (65). Its explicit derivation is in available in S.I. III.
Minimizing Eq. (67) in practice require some approximations. First, the choice of the functionals describing the quantum system and the liquid considerably influences the quality of the predictions. However, determining the appropriate electronic functional for a given solute and the suitable functional for a classical fluid are complex topics that themselves deserve a thorough review. Therefore, we let this discussion aside to focus on the QM/MM coupling and the several strategies adopted in the litterature.
In Petrosyan et al’s joint DFT Petrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007), the interaction of the solvent with both the solute nuclei and electrons is dealt as a whole and the coupling is described by the use of pseudo potential. Thus Eq. (67) solely contains the quantum functional $\mathcal{F}_{\text{qm}}$, the solvent one $\mathcal{F}_{\text{mm}}$ and $\varepsilon_{{\text{qm}}}^{{\text{mm}}}$. The coupling term is modeled as a dielectric screening due to a non local dielectric function depending on the electron density and a mean-field term between both densities interacting through a Gaussian repulsive potential.
In Tang et alTang et al. (2020) Reaction DFT, both electronic and classical densities are not optimized simultaneously. Instead, the quantum functional is optimized in a continuum model. The ground state electronic density is used to compute point charges on the solute atoms which generate an external electrostatic potential in the subsequent classical functional optimization. The interaction between the quantum and the classical parts is similar to Petrosyan’s work, i.e the contribution of the electron and nuclei are gathered together. It is limited to the mean-field term, here a Coulombic term between point charges and a Lennard-Jones term and the correlation functional $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}$ is lacking. Moreover, since both density are not optimized self-consistently the interaction term $\varepsilon_{{\text{qm}}}^{{\text{mm}}}$ only influences the classical optimization and not the quantum one.
In our previous worksJeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024), we also ignored the correlation functional $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}$ in Eq. (65) and limit ourselves to the mean-field term. The interaction potential was also the sum of a Lennard-Jones and a Coulombic term. However, we do not resort to point charges representation to compute the electrostatic energy but directly use the electronic density
$\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ $=\int d{q}\int dr\frac{\rho(q)n_{C}(r)}{|r-{q}|}$ (68)
where $n_{C}$ is the solvent charge density.
We also proposed to couple wave function theory-based methods, such as Hartree-Fock and selected configuration interaction with cDFTLabat, Giner, and Jeanmairet (2024) to approximate the universal Levy-Lieb functional $\mathcal{F}_{\text{qm}}[\rho]$ of Eq. (65).
The results obtained in these worksPetrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007); Tang et al. (2020); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024) clearly indicate that, for ground state properties such as dipoles for instance, the mean-field QM/MM approximation gives already a decent improvement with respect to in vaccuum pure QM calculations.
IV Conclusion
In the present work we have developed exact variational formulations of the Helmholtz free energy of a QM/MM system made of $N_{{\text{qm}}}$ and $N_{{\text{mm}}}$ identical particles. An important result is that the exact Helmholtz free energy of a QM/MM system can be obtained though a minimization of functionals over objects of reduced dimensions, namely the one-body classical and quantum density.
The assumptions of these derivations is a QM/MM partitioning of the system, together with models for both the MM part and the QM/MM interaction. We therefore do not discuss the QM/MM crossover nor how to go beyond the classical models. The starting point of these derivations is the proposal of the QM/MM equivalent to the quantum ground state variational principle or the classical Gibbs measure variational principle (see Sec. II.3). This QM/MM functional is formalized in terms of a $N_{{\text{mm}}}$ classical probability density and an $N_{{\text{qm}}}$ quantum density matrix which also parametrically depends on the $N_{{\text{mm}}}$ classical variables (see Sec. Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ for proofs of the variational property). The latter was proposed thanks to the knowledge of the equilibrium density of the QM/MM system (see Eqs. (14) and (17)) which can be obtained from first principle through a semi classical expansion of a partial Wigner transformation (see Ref. Nielsen, Kapral, and Ciccotti, 2001 and the section II of the supplementary information for a detailed derivation).
Starting from this QM/MM variational principle written in terms of $N$-body objects and thanks to constrained search formalisms similar to that proposed by Levy and Lieb in eDFT, we establish a universal QM/MM functional depending only on the one-body classical and quantum objects (see Sec. III.1). Building on these results, we explicitly make a coupling with the finite temperature eDFT and cDFT functionals, which allows us to rigorously link the two theories for the description of a QM/MM system (see Sec. III.2). We then propose a practical application of this theory in the case of solvation problems, and propose a mean-field treatment. Thanks to these derivations, it becomes clear that the explicit correlation between the classical and quantum electronic systems, were implicitly neglected in our previously reported QM/cDFT calculationsJeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024).
Supplementary materials
The supplementary materials document contains additional information concerning: I) a recall of the variational formulation of statistical quantum mechanics in the canonical ensemble, II) an alternative derivation of the adiabatic QM-MM equilibrium density which does not rely on the Liouville equations or Poisson bracket operators, and III) an extension of the QM/MM functional to the semi-grand-canonical ensemble where the number of classical particles is allowed to vary.
Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$
We wish to show that i) the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ of Eq. (35) is bounded from below by the Helmholtz free energy $F_{0}=-T\log(Z)$ with $Z$ defined in Eq. (17), and ii) that the only minimizer of $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ is $\mathcal{P}_{\text{eq}}$ of in Eq. (14). We will use Jensen’s inequality from convex analysis and probability theory to prove these two statements straightforwardly. This path is similar to that proposed by Friedli and Velenik in the case of classical lattice systems (see Lemma 6.74 in Ref. Friedli and Velenik, 2017) but here applied to the mixed QM/MM case. In order to be able to use Jensen’s inequality, we will first need to show that the couple $(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})$ is indeed a probability measure. While for $\boldsymbol{p}$ it is obvious, it requires a little more attention for $\hat{{\boldsymbol{\rho}}\,}$ as it consists formally in a operator. Then, $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ will be written as an average value on probability measures, and Jensen’s inequality and strict convexity will do the rest.
.1 Rewriting $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ as an average value on probability measures
We use the adiabatic basis ${{\mathcal{B}}({\bf{Q}})=\{{|\Psi_{i}({\bf{Q}})\rangle}},0\leq i\leq\infty\}$ to develop the quantum or mixed quantum-classical operators. More precisely, ${\mathcal{B}}({\bf{Q}})$ consists in the complete set of eigenvectors of the following Schrodinger equation
$\big(\hat{{\bf h}}^{\text{tot}}_{\text{qm}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})\big){|\Psi_{i}({\bf{Q}})\rangle}=E_{i}({\bf{Q}}){|\Psi_{i}({\bf{Q}})\rangle},$ (A.1)
such that the Hamiltonian $\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$ is naturally written as
$\hat{{\bf h}}^{\text{tot}}_{\text{qm}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})=\sum_{i}E_{i}({\bf{Q}}){|\Psi_{i}({\bf{Q}})\rangle}{\langle\Psi_{i}({\bf{Q}})|},$ (A.2)
and the operator $\hat{{\boldsymbol{\rho}}\,}({\bf{Q}})$ is written as follows
$\hat{{\boldsymbol{\rho}}\,}({\bf{Q}})=\sum_{i}{\boldsymbol{\rho}}_{ii}({\bf{Q}}){|\Psi_{i}({\bf{Q}})\rangle}{\langle\Psi_{i}({\bf{Q}})|}.$ (A.3)
Because of the properties of density matrices, the real-numbers ${\boldsymbol{\rho}}_{ii}({\bf{Q}})$ fulfill
${\boldsymbol{\rho}}_{ii}({\bf{Q}})\geq 0,\sum_{i\in{\mathcal{B}}({\bf{Q}})}{\boldsymbol{\rho}}_{ii}({\bf{Q}})=1.$ (A.4)
Therefore, as $\boldsymbol{p}({\bf{Q}})$ is a probability distribution and because of the properties of Eq. (A.4), if we define the function $\mu({\bf{Q}},i)$ as
${\mathbb{R}}^{3N_{{\text{qm}}}}\times{\mathbb{N}}\rightarrow{\mathbb{R}},$ (A.5)
$({\bf{Q}},i)\mapsto\mu({\bf{Q}},i)=\boldsymbol{p}({\bf{Q}}){\boldsymbol{\rho}}_{ii}({\bf{Q}}),$
then $\mu({\bf{Q}},i)$ fulfils the requirements for a probability measure, i.e.
$\mu({\bf{Q}},i)\geq 0\,\,\forall\,\,({\bf{Q}},i),\quad\int\text{d}{\bf{Q}}\sum_{i}\mu({\bf{Q}},i)=1.$ (A.6)
For the sake of compactness, we define ${\mathbf{X}}=({\bf{Q}},i)$, such that we can rewrite the probability measure as
$\mu({\bf{Q}},i)\equiv\mu({\mathbf{X}}),\quad\int\text{d}{\bf{Q}}\sum_{i}\equiv\int\text{d}{\mathbf{X}}.$ (A.7)
The equilibrium density $\hat{\mathcal{P}}_{\text{eq}}$ of Eq. (14) can be written using the adiabatic basis of Eq. (A.2) as follows
${\hat{\mathcal{P}}_{\text{eq}}}({\bf{Q}})$ $=\frac{1}{Z}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}e^{-\beta\hat{{\bf V}}^{\text{tot}}_{\text{mm}}({\bf{Q}})}$ (A.8)
$=\frac{1}{Z}\sum_{i}e^{-\beta\epsilon({\bf{Q}},i)}{|\Psi_{i}({\bf{Q}})\rangle}{\langle\Psi_{i}({\bf{Q}})|},$
which corresponds to the measure $\mu({\mathbf{X}})=\mu_{\text{eq}}({\mathbf{X}})$ defined as follows
$\mu_{\text{eq}}({\mathbf{X}})\equiv\frac{e^{-\beta\epsilon({\mathbf{X}})}}{Z},$ (A.9)
where
$\epsilon({\mathbf{X}})=\epsilon({\bf{Q}},i)\equiv E_{i}({\bf{Q}})+\hat{{\bf V}}^{\text{tot}}_{\text{mm}}({\bf{Q}}).$ (A.10)
We introduce the notation for an average value over a probability measure $\mu({\mathbf{X}})$ as
$\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})f({\mathbf{X}})=\langle f\rangle_{\mu},$ (A.11)
such that the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ can be rewritten as an average value on the measure $\mu({\mathbf{X}})$
$F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]=$ $\text{Tr}\big\{\hat{{\boldsymbol{\rho}}\,}\boldsymbol{p}\big(\hat{{\bf V}}^{\text{tot}}_{\text{mm}}+\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}+T\log(\hat{{\boldsymbol{\rho}}\,}\boldsymbol{p})\big)\}$ (A.12)
$=$ $\int\text{d}{\bf{Q}}\sum_{i}\mu({\bf{Q}},i)\big(\epsilon({\bf{Q}},i)+T\log(\mu({\bf{Q}},i))$
$=$ $\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\big(\epsilon({\mathbf{X}})+T\log(\mu({\mathbf{X}}))$
$=$ $-T\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\log(\frac{e^{-\beta\epsilon({\mathbf{X}})}}{\mu({\mathbf{X}})})$
$=$ $-T\langle\log(\frac{e^{-\beta\epsilon}}{\mu})\rangle_{\mu}$
$\equiv$ $F[\mu].$
According to Eq. (A.12), the evaluation $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ at $\hat{\mathcal{P}}_{\text{eq}}$ leads straightforwardly to
$F[\mu_{\text{eq}}]=$ $-T\log(Z),$ (A.13)
which is of course the correct Helmholtz free energy. Then two questions remain: i) is $T\log(Z)$ a global minima and ii) is $\hat{\mathcal{P}_{\text{eq}}}$ the only minimizer ?
.2 Use of Jensen inequality to prove variational character
As the function $\log(x)$ is strictly convex, we will now use Jensen’s inequality from convex analysis and probability measures to prove that $F[\mu]$ is bounded from below by the Helmholtz free energy and that $\mu_{\text{eq}}$ is the only minimizer. The Jensen’s inequality can be stated as follows.
If $\varphi$ is a convex function (i.e. always above its tangent lines), $\mu({\mathbf{X}})$ a probability measure on ${\mathbf{X}}\in E$ where $E$ is a normed vector space, $g({\mathbf{X}})$ a function from $E$ to $\mathbb{R}$}, then Jensen’s equality states that
$\varphi(\langle g\rangle_{\mu})\leq\langle\varphi(g)\rangle_{\mu},$ (A.14)
or written in terms of integrals
$\varphi\bigg(\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})g({\mathbf{X}})\bigg)\leq\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\varphi\big(g({\mathbf{X}})\big).$ (A.15)
As $\log(x)$ is convex, the Jensen’s equality can be applied to $F[\mu]$ of Eq. (A.12) by setting
$g({\mathbf{X}})=\frac{e^{-\beta\epsilon({\mathbf{X}})}}{\mu({\mathbf{X}})},$ (A.16)
$\varphi(x)=-T\log(x).$
Therefore the inequality of Eq. (A.14) becomes
$T\log\langle\frac{e^{-\beta\epsilon}}{\mu}\rangle_{\mu}\leq-T\langle\log(\frac{e^{-\beta\epsilon}}{\mu})\rangle_{\mu},$ (A.17)
and as
$\langle\frac{e^{-\beta\epsilon}}{\mu}\rangle_{\mu}$ $=\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\frac{e^{-\beta\epsilon(x))}}{\mu({\mathbf{X}})}$ (A.18)
$=Z,$
one can rewrite Eq. (A.17) as
$-T\log(Z)\leq F[\mu],$ (A.19)
which guarantees the variational character. Eventually, the strict convexity of the $\log(x)$ function implies that the Jensen’s inequality becomes an equality if and only if $g({\mathbf{X}})=cst$, which translates into
$g({\mathbf{X}})=cst\Leftrightarrow\frac{e^{-\beta\epsilon({\mathbf{X}})}}{\mu({\mathbf{X}})}=cst\Leftrightarrow\mu({\mathbf{X}})\propto e^{-\beta\epsilon({\mathbf{X}})},$ (A.20)
therefore implying that $\mu_{\text{eq}}({\mathbf{X}})$ is the only minimizer. We point out that this proof essentially consists in the Gibb’s inequality re-framed with convex analysis vocabulary.